### v1

# Fibred surfaces and their unitary rank

Abstract
Let $f\colon S\to B$ a fibred surface with fibres of genus $g\geq 2$ and base curve of genus $b$. Let $u_{f}$ be its unitary rank. We prove many new slope inequalities involving $u_{f}$ and some other invariants of the fibration. As applications:
(1) we prove a new Xiao-type bound on $u_{f}$ with respect to $g$ for non-isotrivial fibrations:
$u_{f}<g\frac{5g-2}{6g-3}.$
In particular this imples that if $f$ is not locally trivial and $u_{f}=g-1$ is maximal, then $g\leq 6$;
(2) we prove a result in the direction of the Coleman-Oort conjecture: a new constrain on the rank of the (-1,0) part of the maximal unitary Higgs subbundle of a curve generically contained in the Torelli locus;
(3) we study in the detail the extremal case (for non-isotrivial $f$) where $u_{f}=g-1$, giving many constrains for the case $g=6$, $u_{f}=5$.
2010 Mathematics Subject Classification. 14J10, 14D06, 14D07, 14G35, 14H40. Key words and phrases: fibred varieties, slope inequality, families of Jacobians, Hodge bundle, Coleman-Oort conjecture
1 Introduction
This paper deals with the geography of fibred surfaces and some applications. The main results can be divided in four sections, which we list in order of appearance.
New slope inequalities
Let $f\colon S\to B$ a relatively minimal fibred surface. We prove new slope inequalities, i.e. bounds of the form
$K_{f}^{2}\geq\alpha\chi_{f},$
where $\alpha$ is a function of some invariants of the fibration. In particular, in our $\alpha$ will be a function depending on the genus $g$ of the general fibre, on the unitary rank $u_{f}$, and in some cases on the Clifford index of the general fibre $c_{f}$, and on the rank $r_{1}$ of the first subbundle of the Harder-Narasimhan filtration of the Hodge bundle $f_{*}\omega_{f}$ (see Section 2 for the definitions). Some of the most significant inequalities we prove are the following:
Theorem 1.1 (Theorem 4.1, Theorem 4.25).
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$.
  • The following inequality holds:
    $K_{f}^{2}\geq\frac{4g(g-1)}{(2g-1)(g-u_{f})}\chi_{f}.$ (1.2)
    Moreover, equality holds in (1.2) if and only if $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$, where $\Gamma$ is a section of $f$.
  • Assume that the ample summand $\mathcal{A}$ in the second Fujita decomposition (2.7) of $f_{*}\omega_{f}$ is semistable. Then the following inequality holds:
    $K_{f}^{2}\geq\left(\frac{4(g-1)-2u_{f}}{(g-u_{f})}+\frac{2u_{f}}{(2u_{f}+1)(g-u_{f})}\right)\chi_{f}.$ (1.3)
    Moreover, if equality holds in (1.3), then
    • if $u_{f}\leq g-2$ then $F$ is hyperelliptic;
    • if $u_{f}=g-1$ then $F$ is trigonal.
  • Assume that $\mathcal{A}$ is semistable and that $u_{f}\leq c_{f}$. Then the following inequality holds:
    $K_{f}^{2}\geq\left(\frac{4(g-1)-u_{f}}{(g-u_{f})}+\frac{u_{f}}{(u_{f}+1)(g-u_{f})}\right)\chi_{f}.$ (1.4)
Remark 1.5.
The first inequality (1.2) is an extension of a result of Konno in [34], where he proved the same inequality with the relative irregularity $q_{f}$ instead of $u_{f}$. The second and third inequalities are a sharpening of an inequality due to the author with Barja in [7] and of inequalities obtained by the author with Riva in [48].
A new Xiao type inequality for the unitary rank
Using inequality (4.3) we extend to the unitary rank the bound obtained by Konno in [34] that was itself a sharpening of the famous Xiao’s bound [52]. In particular, we obtain that:
Theorem 1.6 (Theorem 6.1, Proposition 6.4).
For a relatively minimal non-isotrivial fibred surface of genus $g$ and unitary rank $u_{f}$, it always holds
$u_{f}<g\frac{5g-2}{6g-3}.$ (1.7)
Moreover, the following hold:
  1. 1.
    if the unitary rank is maximal $u_{f}=g-1$, then necessarily $g\leq 6$;
  2. 2.
    if the unitary rank $u_{f}=g-2$, then necessarily $g\leq 11$.
This is, up to now, the sharpest known bound for $u_{f}$; see Section 2 for the known results and conjectures.
Results on the case with high unitary rank
Recall that Pirola proved in [47] that for a relatively minimal fibration such that $u_{f}=g-1$, then up to a base change $q_{f}=u_{f}$. Also the following result is essentially due to Pirola (in a private communication with the author) and to the already known results [6], [30].
Theorem 1.8 (Proposito 5.1, Theorem 5.2).
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then $f$ is a non-isotrivial trigonal fibration, the smooth fibres $F_{t}$ are a covering of a non-isotrivial family of elliptic curves. Moreover, $f$ is not not Kodaira, i.e. $K_{f}^{2}<12\chi_{f}$.
Applications to the Coleman-Oort conjecture
Having proved these sharper inequalities, we can apply them to the techniques developed by Viehweg and Zuo and then by Chen, Lu and Zuo (see [50], [17], [40]) and give a result that goes in the direction of the Coleman-Oort conjecture on the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties of genus $g$ (see Section 7 for the definitions and some explanation):
Theorem 1.9 (Theorem 7.1).
Let $C\subset\mathcal{A}_{g}$ be a curve with Higgs bundle decomposition $\mathcal{E}_{C}=\mathcal{A}_{C}\oplus\mathcal{U}_{C}$, where $\mathcal{U}_{C}$ is the maximal unitary Higgs subbundle.
If $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}\geq g\frac{5g-2}{6g-3}$, then $C$ is not contained generically in the Torelli locus.
Moreover, if $C$ is a Shimura curve such that $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}\geq\frac{4g-1+\sqrt{16g^{2}-36g+21}}{10}$ then $C$ is not contained generically in the Torelli locus.
This is a sharpening of a theorem of Chen, Lu and Zuo [17, Thm 1.1.2]: see Remark 7.5.
Results on the case with $g=6$, $u_{f}=5$
Eventually we can use our arguments, together with some other result on trigonal curves, to give some information on the extremal cases. First recall that Beorchia, Pirola and Zucconi proved in [13] that if $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $6$ and with $q_{f}=5$ then $F$ is trigonal of special Maroni invariant.
Theorem 1.10 (Proposition 8.1, Corollary 8.8).
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $6$ and with $u_{f}=5$. Then:
  1. 1.
    the surface $S$ has positive index $\tau(S)>0$.
  2. 2.
    The schematic base locus associated to the morphism $f^{*}\mathcal{A}\to\omega_{f}$ has a horizontal component $G$ of multiplicity at least 5.
  3. 3.
    If the multiplicity of $G$ is at least $6$ then $G$ is a section and the point $G\cap F$ on the general fibre is a Weierstrass point and we can list its possible gap sequences.
  4. 4.
    We can track down the possibilities for the numerical class of $K_{f}$, reducing them to the following cases:
    • (a)
      $K_{f}\equiv 10\Gamma+\mu_{1}F$;
    • (b)
      $K_{f}\equiv 9\Gamma+\Gamma^{\prime}+\mu_{1}F$;
    • (c)
      $K_{f}\equiv 6\Gamma+4\Gamma^{\prime}+\mu_{1}F$;
    • (d)
      • either $K_{f}\equiv 5\Gamma+5\Gamma^{\prime}+\mu_{1}F$,
      • or $K_{f}\equiv 5\Sigma$
    where $\Gamma$ and $\Gamma^{\prime}$ are distinct sections of $f$ and $\Sigma$ is a bisection of $f$.
Remark 1.11.
So we have given many constrains to the case of non locally trivial fibrations with $g=6$, $u_{f}=5$, which add up to the ones already given by Beorchia, Pirola, Zucconi [13], and also to the explicit conditions found by Colombo, Frediani and Pirola in the recent paper [19, Section 10]. Still, the existence of such fibred surfaces is not known, the conjecture being that they do not exist. On the other hand, for $g\leq 4$ these fibration exists -the first example being given by Pirola in [46]- and have been thoroughly investigated ([1], [44], [18]).
Moreover, we can find some conditions for the case $g=5$ and $u_{f}=4$, listed in Proposition 8.17.
For these last results we develop a finer slope inequality, applicable to these extremal cases (Proposition 6.4).
A few words about the proofs. We work in the framework of the celebrated Xiao’s method; however, a new key ingredient is a re-elaboration of a result due to Konno in [34], where he gives an original construction of a nef divisor on the surface. Moreover, we use in many points the algebraic index theorem, in particular to give characterizations of the extremal cases. In the last section, with an ad-hoc argument, we prove a very strong slope inequality holding in the extremal cases $u_{f}=g-1$. This, together with the use of results of Coppens ([20, 21, 22]) and Brundu-Sacchiero ([14]) allows us to give the results about this extremal case.
Acknowledgements This article would not have been written without the encouragement and stimulation of Gian Pietro Pirola, whom I thank heartily.
2 Fibred surfaces and their geography
We call fibred surface or sometimes simply fibration the data of a morphism $f\colon S\rightarrow B$ from a smooth projective surface $S$ to a smooth projective curve $B$ which is surjective with connected fibres. We denote with $b=g(B)$ the genus of the base curve. A general fibre $F$ is a smooth curve and its genus $g=g(F)$ is by definition the genus of the fibration. From now on, we consider fibrations of genus $g\geq 2$.
We say that $f$ is relatively minimal if it does not contain any $(-1)$-curves in its fibres. This condition is equivalent to $K_{f}$ being a relatively nef divisor. It indeed holds that $K_{f}$ is nef ,as proved for instance in [5]. Given any fibration of genus $g\geq 2$ we can contract all the $(-1)$-curves contained in the fibres and obtain a unique relatively minimal genus $g$ fibration $\overline{f}\colon\overline{S}\to B$. We will always assume, unless differently stated, that a fibred surface is relatively minimal.
Given a fibred surface, it is natural to define some relative numerical invariants, as follows:
  • $K_{f}^{2}=K_{S}^{2}-8(g-1)(b-1)$ the self-intersection of the relative canonical divisor;
  • $\chi_{f}:=\chi(\mathcal{O}_{S})-(g-1)(b-1)$ the relative Euler characteristic;
  • $e_{f}:=e(S)-e(B)e(F)=e(S)-4(g-1)(b-1)$ the relative topological characteristic (with $e(X)$ topological characteristic of $X$);
  • $q_{f}:=q-b$ the relative irregularity, with $q=h^{1}(S,\mathcal{O}_{S})$ irregularity of $S$ (see also Remark 2.8);
  • We can also consider as invariants of the fibration the gonality and the Clifford index of $F$ which we denote $gon(F)$ and $c_{f}=\mbox{{Cliff}}(F)$ respectively.
From Groethendieck-Riemann-Roch theorem we have Noether’s relation
$12\chi_{f}=K_{f}^{2}+e_{f}.$ (2.1)
We say that a fibred surface is: smooth if every fibre is smooth; isotrivial if all smooth fibres are mutually isomorphic; locally trivial if $f$ is smooth and isotrivial (equivalently $f$ is a fibre bundle by the Grauert-Fisher Theorem [9, Theorem I.10.1]); trivial if $S$ is birationally equivalent to $F\times B$ and $f$ corresponds to the projection on $B$. If $b>0$ and $f$ is relatively minimal this is equivalent to $S=F\times B$. Eventually, we say that $f$ is a Kodaira fibration if it is smooth and non isotrivial.
Let us first summarize what we know about the non-negativity of these invariants.
Theorem 2.2.
Let $f\colon S\to B$ a relatively minimal fibred surface. The following results hold:
  1. (1)
    $K_{f}^{2}\geq 0$ and $K_{f}^{2}=0$ if and only if $f$ is locally trivial (see [5] and Remark 2.16);
  2. (2)
    $\chi_{f}\geq 0$ and $\chi_{f}=0$ if and only if $f$ is locally trivial (see [9]);
  3. (3)
    $e_{f}\geq 0$ and $e_{f}=0$ if and only if $f$ is smooth (see [11]);
  4. (4)
    $q_{f}\leq g$ and equality holds if and only if $f$ is trivial (see [12] and also Remark 2.9).
Definition 2.3.
The rank $g$ vector bundle $f_{*}\omega_{f}$ is called the Hodge bundle of the fibred surface.
By using Grothendieck-Riemann-Roch theorem or Leray’s spectral sequence, we see that $\deg f_{*}\omega_{f}=\chi_{f}$ for any fibration $f$.
Definition 2.4.
A vector bundle $\mathcal{E}$ over a smooth curve is called nef if one of the following equivalent properties holds: (1) the corresponding tautological sheaf $\mathcal{O}_{\mathbb{P}(\mathcal{E})}(1)$ is nef; (2) if all its quotients have non-negative degree.
Let us recall some important results on the Hodge bundle.
Theorem 2.5.
The Hodge bundle of a fibration $f\colon S\to B$ of genus $g\geq 1$ is a rank $g$ nef vector bundle over $B$ ([33]).
The Hodge bundle can be decomposed in two ways as a direct summand of vector sub-bundles as follows:
  • (First Fujita decomposition [25])
    $f_{*}\omega_{f}=\mathcal{O}_{B}^{\oplus q_{f}}\oplus\mathcal{G},$ (2.6)
    where $\mathcal{G}$ is nef and $H^{0}(B,\mathcal{G}^{\vee})=H^{1}(B,\mathcal{G}\otimes\omega_{B})=0$;
  • (Second Fujita decomposition [26] [16])
    $f_{*}\omega_{f}=\mathcal{A}\oplus\mathcal{U},$ (2.7)
    where $\mathcal{A}$ ample and $\mathcal{U}$ unitary flat.
Remark 2.8.
So, we see that $q_{f}$ is the rank of the biggest trivial subbundle of $\mathcal{E}$, moreover, we see that $\chi_{f}=\deg\mathcal{G}=\deg\mathcal{A}$. We can also see directly from the first decomposition that $q_{f}\leq g$ and that if $q_{f}=g$ then $\chi_{f}=0$, and so $f$ is locally trivial (but Beauville’s result (4) of Theorem 2.2 is stronger).
Following [30], we define the unitary rank $u_{f}$ of the fibred surface to be the rank of $\mathcal{U}$.
Remark 2.9.
Comparing the two decompositions, since every trivial bundle is unitary flat, we have that:
$\mathcal{O}_{B}^{\oplus q_{f}}\subseteq\mathcal{U},$
and then it holds that $q_{f}\leq u_{f}$. Observe that $\deg\mathcal{U}=0$, so we have that $\chi_{f}=\deg\mathcal{A}$, and moreover, $\mathcal{U}$ is the biggest degree 00 sub-bundle of $f_{*}\omega_{f}$, so $u_{f}=g$ if and only if $\chi_{f}=0$ (equivalently $f$ is locally trivial): compare with (2) of Theorem 2.2.
One of the most important areas of study in this field is the so called geography of fibred surfaces, i.e. the study of the inequalities holding between the invariants. We will list three kinds of these inequalities. For very nice expositions of these kind of problems see also [42] and [10].
Inequalities between the invariants: bounds on the ratio $K_{f}^{2}/\chi_{f}$
First of all, from Noether’s relation (2.1) and ((3)) of Theorem 2.2 we have:
Theorem 2.10.
For any fibred surface of genus $g\geq 2$, we have
$K_{f}^{2}\leq 12\chi_{f}.$ (2.11)
Moreover, if $f$ is relatively minimal and equality holds, then $f$ is a Kodaira fibration.
Remark 2.12.
Given any relatively minimal non-locally trivial fibred surface $f\colon S\to B$ over a base of genus $b\geq 2$, the surface $S$ is minimal and of general type. So, Bogomolov-Miyaoka-Yau inequality holds: $K_{S}^{2}\leq 9\chi(\mathcal{O}_{S})$. This on the relative invariants implies that
$K_{f}^{2}\leq 9\chi_{f}+(g-1)(b-1),$
which is not a “slope inequality” in the sense made explicit in the introduction, nevertheless it gives interesting informations.
Remark 2.13.
Note moreover that the ratio 8 is particularly significant for the geometry of the surface. Indeed, consider the topological index $\tau(S)$ of the surface, i.e. the index of the intersection form on $H^{2}(S,\mathbb{C})$ (the difference between the number of its positive and negative eigenvalues). We have that $\tau(S)=\frac{1}{3}(K_{S}^{2}-2c_{2}(S))$. So,
$K_{f}^{2}-8\chi_{f}=K_{S}^{2}-8\chi(\mathcal{O}_{S})=K_{S}^{2}-2c_{2}(S)=3\tau(S).$
For what concerns lower bounds on the ratio $K_{f}^{2}/\chi_{f}$, the first and most famous result is the so-called canonical slope inequality:
Theorem 2.14.
[Xiao [51], Cornalba-Harris [23, Sto], see also [4, Thm. 4.5], Konno [35]]
Let $f\colon S\to$ be a relatively minimal fibred surface, then the following inequality holds:
$K_{f}^{2}\geq\frac{4g-4}{g}\chi_{f}.$ (2.15)
If $f$ is non-locally trivial and the equality in (2.15) is reached, then $F$ is hyperelliptic and $q_{f}=0$.
Remark 2.16.
Suppose that $K_{f}^{2}=0$. Then by the slope inequality we have $\chi_{f}=0$ so $f$ is locally trivial. If, on the other hand, $f$ is locally trivial; then by $(ii)$ we have $\chi_{f}=0$, then by Noether’s relation and the non-negativity of $e_{f}$ we have $K_{f}^{2}=0$.
Remark 2.17.
Even from the classical results of Theorem 2.14, it is natural to ask for bound increasing with $c_{f}$ and with $q_{f}$. Many authors have contributed to this topic, here we just recall [36], [7], [8], [37], [39].
Inequalities between the invariants: Xiao’s type of bounds
As mentioned in $q_{f}\leq g$, and equality holds if and only if $f$ is trivial. So, it is natural to ask: if $f$ is not trivial, what is the inequality between $g$ and $q_{f}$? From now on we assume $f$ is non-trivial. Xiao initiated the study in [52] by proving that if $b=0$
$q_{f}\leq\frac{g+1}{2}.$ (2.18)
For arbitrary $b$ he proved the bound
$q_{f}\leq\frac{5g+1}{6}.$ (2.19)
He also conjectured that the bound (2.18) holds for any non-trivial fibration. Serrano [49] proved that if $f$ is isotrivial (but not trivial), then (2.18) holds. Cai in [15] proved that if $f$ is non-isotrivial and the general fibre is either hyperelliptic or bielliptic, the same bound holds. In the article that inspired our present work Konno improved the bound (2.18) to:
$q_{f}\leq g\frac{5g-2}{6g-3}$ (2.20)
Pirola gave a counterexample of genus $3$ to Xiao’s conjecture in [46], by constructing for non-isotrivial fibrations a higher Abel-Jacobi map. More counterexamples have been found later by Albano and Pirola in [1].
In [6] Barja, González and Naranjo proved the following: if $f$ is non-isotrivial, then $q_{f}\leq g-c_{f}$; moreover, they made the following modified conjecture:
Conjecture 2.21 (Barja, González, Naranjo; Modified Xiao’s conjecture for the relative irregularity).
For any non-isotrivial fibred surface $f\colon S\to B$ of genus $g\geq 2$ it holds
$q_{f}\leq\left\lceil\frac{g+1}{2}\right\rceil.$ (2.22)
Favale-Naranjo-Pirola [24] have proven the stronger inequality $q_{f}\leq g-c_{f}-1$ for families of plane curves of degree $d\geq 5$.
It is natural to ask wether these bounds can be extended to the unitary rank $u_{f}$.
In the paper [30] the author with González and Torelli proved the inequality $u_{f}\leq g-c_{f}.$ Moreover, if the general fibre is a plane curve of degree $d\geq 5$, then $u_{f}\leq g-c_{f}-1.$
Remark 2.23.
If $\mathcal{U}$ has finite monodromy, then up to a base change it becomes trivial, and so coincides with $q_{f}$. However, the unitary summand $\mathcal{U}$ of the Hodge bundle can have infinite monodromy as proved by Catanese and Dettweiler in [16]. Moreover, modified Xiao’s conjecture does not hold for $u_{f}$: this follows from [16, Remark 38 ] and from the construction of Lu [41]. See also [30, Conjecture 1.1].
3 Main technical results
3.1 The Harder-Narashiman sequence of a vector bundle
Let us briefly recall the notion of the Harder-Narasimhan sequence of a vector bundle $\mathcal{E}$ over a curve $B$ (see [31]). It is the unique filtration of subbundles
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\ldots\subset\mathcal{E}_{l}=\mathcal{E}$
satisfying the following assumptions:
  • for any $i=0,\ldots l$ $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is a $\mu$-semistable vector bundle;
  • if we set $\mu_{i}:=\mu(\mathcal{E}_{i}/\mathcal{E}_{i-1})$, we have that $\mu_{i}>\mu_{i-1}$.
Note that $\mu_{1}>\mu(\mathcal{E})>\mu_{l}$, unless $\mathcal{E}$ is $\mu$-semistable, in which case $1=l$ and $\mu(\mathcal{E})=\mu_{1}$. Note moreover that, setting $r_{0}:=0$, we can express the degree of $\mathcal{E}$ as a combinations of the $\mu_{i}$’s and $r_{i}$’s:
$\deg(\mathcal{E})=\sum_{i=1}^{l}\mu_{i}(r_{i}-r_{i-1}).$ (3.1)
Let us call $H$, respectively $\Sigma$ a tautological divisor, respectively a fibre on the projective bundle $\pi\colon\mathbb{P}_{B}(\mathcal{E})\to B$. Recall that $\mbox{{Pic}}(\mathbb{P}_{B}(\mathcal{E}))=\pi^{*}\mbox{{Pic}}(B)\oplus\mathbb{Z}[H]$. The following result of Miyaoka-Nakayama [43, 45] determines the form of the nef and pseudoeffective cones of $\mathbb{P}_{B}(\mathcal{E})$.
Theorem 3.2.
[Miyaoka-Nakayama]
With the above notations, given $k\in\mathbb{Z}$ and $D$ a divisor on $B$,
  1. 1.
    The divisor $kH-\pi^{*}D$ is pseudo-effective (i.e. it is a limit of effective divisors) if and only if $k\geq 0$ and $\deg D\leq k\mu_{1}$.
  2. 2.
    The divisor $kH-\pi^{*}D$ is nef if and only if $k\geq 0$ and $\deg D\leq k\mu_{l}$.
Remark 3.3.
Although the theorem is stated for divisors, it can be readily extended to $\mathbb{Q}$-divisors, and it is in this form that we will use it.
Remark 3.4.
From Theorem 3.2 (or from the second definition of nefness in Definition 2.4) we see that $\mathcal{E}$ is nef if and only if the smallest slope $\mu_{l}$ is greater or equal to 00. In this case, setting $\mu_{l+1}:=0$, we can reformulate equation (3.1) as follows (see [48, Rem. 3.15]):
$\deg(\mathcal{E})=\sum_{i=1}^{l}r_{i}(\mu_{i}-\mu_{i+1}).$
3.2 The Harder-Narashiman sequence of the Hodge bundle
Let us consider the case of a fibred surface $f\colon S\to B$, and consider its Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$. As $\mathcal{E}$ is nef by Theorem 2.5, we have that $\mu_{l}$ is greater or equal to 00.
Remark 3.5.
Observe that the unitary flat summand $\mathcal{U}$ in (2.7) is the biggest subbundle of $\mathcal{E}$ with degree 00. So, in particular we have that $\mu_{l}=0$ if and only if $\mathcal{U}\not=0$ if and only if $u_{f}\not=0$.
Moreover, if $\mu_{l}=0$, then $\mathcal{E}_{l-1}$ is precisely $\mathcal{A}$.
Remark 3.6.
Observe moreover that we have in general:
$\mu_{l}\leq\frac{\chi_{f}}{g-u_{f}}\leq\frac{\chi_{f}}{g-q_{f}}\leq\frac{\chi_{f}}{g}\leq\mu_{1}$
and
  • (i)
    $\mu_{1}=\chi_{f}/g=\mu(f_{*}\omega_{f})$ if and only if $f_{*}\omega_{f}=\mathcal{A}$ is semistable.
  • (ii)
    $\mu_{1}=\chi_{f}/(g-u_{f})=\mu(\mathcal{A})$ if and only if $\mathcal{A}$ is semistable.
  • (iii)
    $\mu_{1}=\chi_{f}/(g-q_{f})=\mu(\mathcal{G})$ if and only if $u_{f}=q_{f}$ and $\mathcal{A}$ is semistable.
3.3 Xiao’s set-up
We follow the set-up given by Xiao in its seminal paper [51]. Let us consider the Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$, and its Harder-Narasimhan sequence as in subsection 3.2. For any $i=1,\ldots l$, we have that the sheaf homomorphism
$f^{*}\mathcal{E}_{i}\longrightarrow f^{*}f_{*}\omega_{f}\longrightarrow\omega_{f}$
induces a rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$ (see [32, Pro. 7.12]).
Let $H_{i}$ be the tautological divisor on $\mathbb{P}_{B}(\mathcal{E}_{i})$, let $M_{i}:=\phi_{i}^{*}(H_{i})$ on $S$.
Remark 3.7.
By Theorem 3.2, observing that $\mu_{i}$ is the final slope of $\mathcal{E}_{i}$ for all $i$, we have that $H_{i}-\mu_{i}\Sigma_{i}$ is a nef $\mathbb{Q}$-divisor and $H_{i}-\mu_{1}\Sigma_{i}$ is a pseudoeffective $\mathbb{Q}$-divisor on $\mathbb{P}_{B}(\mathcal{E}_{i})$.
Hence, $\phi_{i}^{*}(H_{i}-\mu_{i}\Sigma_{i})=M_{i}-\mu_{i}F$ is a nef $\mathbb{Q}$-divisor on $S$ and $\phi_{i}^{*}(H_{i}-\mu_{1}\Sigma_{i})=M_{i}-\mu_{1}F$ is a pseudoeffective $\mathbb{Q}$-divisor on $S$.
Moreover, for any $i=1,\ldots l$, call $Z_{i}$ the divisorial base locus of $\phi_{i}$, i.e. the effective divisor in $S$ such that $\mathcal{E}_{i}\subseteq f_{*}\omega_{f}(-Z_{i})$ and such that the evaluation homomorphism
$f^{*}\mathcal{E}_{i}\longrightarrow f^{*}f_{*}\omega_{f}(-Z_{i})\longrightarrow\omega_{f}(-Z_{i})$
is generically surjective. Of course
$Z_{1}\geq Z_{2}\geq\ldots\geq Z_{l}=0.$
Due to the importance of this result in what follows, we give the proof of this lemma:
Lemma 3.8.
With the above notations, for any $i=1,\ldots l$, the divisor $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$,
Proof.
The sheaf homomorphism $f^{*}\mathcal{E}_{i}\to f^{*}f_{*}\omega_{f}(-Z_{i})\to\omega_{f}(-Z_{i})$ is surjective in codimension 2. Let us consider a chain of blow ups $\rho\colon\widetilde{S}\to S$ resolving the points of indeterminacy of the associated map $\phi_{i}$: there exists a morphism $\tilde{\phi}_{i}\colon\widetilde{S}\to\mathbb{P}_{B}(\mathcal{E}_{i})$ such that $\tilde{\phi}_{i}=\phi_{i}\circ\rho$. We thus have the following commutative diagram:
[…]
Let $\tilde{f}=f\circ\rho\colon\widetilde{S}\to B$ be the associated fibration; note that it is non relatively minimal unless $\phi_{i}$ already is a morphism. Let $E_{i}$ be the exceptional divisor of $\rho$.
The morphism $\tilde{\phi}_{i}$ is associated to the line bundle $\mathcal{L}=\rho^{*}\omega_{f}(-Z_{i})\otimes\mathcal{O}_{\widetilde{S}}(-E_{i})$ on $\widetilde{S}$, and we have the surjective sheaf homomorphism ([32, Ch II, Prop 7.11]):
$\tilde{f}^{*}\mathcal{E}_{i}\longrightarrow{\tilde{f}}^{*}{\tilde{f}}_{*}\mathcal{L}\longrightarrow\mathcal{L}.$
So, from [32, Prop 7.12] we have that $\mathcal{L}\cong\widetilde{\phi_{i}}^{*}(\mathcal{O}_{\mathbb{P}_{B}(\mathcal{E}_{i})}(H_{i}))$. Now, given any irreducible curve $C\subset S$, we have by using the push-pull formula that
$\begin{split}(M_{i}\cdot C)&=(\phi_{i}^{*}(H_{i})\cdot C)=(\tilde{\phi}_{i}^{*}(H_{i})\cdot\rho^{*}(C))=(\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C))+(E_{i}\cdot\rho^{*}(C))=\\ &=(\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C))=(K_{f}-Z_{i}\cdot C),\\ \end{split}$
as wanted. ∎
As a direct consequence of the previous discussion and of Lemma 3.8 we have the following:
Proposition 3.9.
[Xiao] With the above notations, we have that for any $i=1,\ldots l$:
  • (i)
    $K_{f}-Z_{i}-\mu_{i}F$ is nef;
  • (ii)
    $K_{f}-Z_{1}-\mu_{1}F$ is pseudoeffective.
Remark 3.10.
Note that the divisors $Z_{i}$’s can have vertical components. In case they have fibres as components, say $Z_{i}=bF+Z^{\prime}_{i}$ with $b\geq 0$ and the support of $Z_{i}^{\prime}$ containing no fibre, then we can observe the following:
  • (i)’
    $K_{f}-Z^{\prime}_{i}-\mu_{i}F$ is nef;
  • (ii)’
    $K_{f}-Z^{\prime}_{1}-\mu_{1}F$ is pseudoeffective.
Indeed, for (i)’, let $C$ be any irreducible curve in $S$. If $C$ is vertical then $C\cdot F=0$ and so
$(K_{f}-Z^{\prime}_{i}-\mu_{i}F)\cdot C=(K_{f}-Z_{i}-\mu_{i}F)\cdot C\geq 0.$
In case $C$ is horizontal then $C\cdot F>0$, so
$(K_{f}-Z^{\prime}_{i}-\mu_{i}F)\cdot C=(K_{f}-Z_{i}-\mu_{i}F)\cdot C+bF\cdot C\geq 0.$
For (ii)’ just observe that $K_{f}-Z^{\prime}_{1}-\mu_{1}F=K_{f}-Z_{1}-\mu_{1}F+bF$, so it is certainly pseudoeffective. For this reason in the following we can assume that the $Z_{i}$’s can only have vertical components that are not fibres. We can not suppose that $Z_{i}$ has no vertical components, because (i)’ is not guaranteed if we eliminate those components.
Another approach to get rid of fibres in the base divisors -the one used by Xiao in [51]- is to tensor $\mathcal{E}_{i}$ with a sufficiently ample line bundle $\mathcal{B}$ on $B$ and to consider $L$ the linear subsystem of $|\omega_{f}\otimes f^{*}\mathcal{B}|$ corresponding to sections in $H^{0}(B,\mathcal{E}_{i}\otimes\mathcal{B})$. Then one defines $Z_{i}$ to be the fixed part of $L$ and verifies that this construction is independent on the choice of the ample line bundle $\mathcal{B}$.
Notation 3.11.
Let us call $d_{i}=M_{i}\cdot F$, the degree of the linear series $|V_{i}|$ induced by $M_{i}$ on the general fibre, and let $a_{i}=(Z_{i}\cdot F)=2g-2-d_{i}$. The divisor ${Z_{i}}_{|F}$ on $F$ is the base locus of the linear system $|V_{i}|$ seen as a subcanonical system.
Also call $r_{i}:=\mbox{{rank}}\mathcal{E}_{i}$, so that $|V_{i}|$ is associated to a base point free subcanonical $g^{d_{i}}_{r_{i}-1}$ on $F$.
Remark 3.12.
Observe that $d_{1}\leq d_{2}\leq\ldots d_{l}=2g-2$, and also note that $d_{i}>0$ (i.e. $a_{i}<2g-2$) if $i>1$, because if $i>1$ the series $|V_{i}|$ has dimension $r_{i}-1>0$. Moreover, $d_{1}=0$ if and only if $r_{1}=1$.
3.4 Konno’s idea: a new nef divisor on $S$
Till now we just introduced the framework of Xiao’s technique. Now we turn to an original idea of Konno, which consists in relating the integers $a_{i}$ with the multiplicity of the irreducible components of $Z_{i}$, and then in using the Hurwitz formula to produce new nef divisors on $S$.
Proposition 3.13 (Konno, [34]).
With the above notation, let $Z_{i}=\sum_{j}m_{j}G_{j}$ be the irreducible decomposition of $Z_{i}$, and let $\alpha_{i}:=\max_{j}\{m_{j}\mid F\cdot G_{j}>0\}$. Then the $\mathbb{Q}$-divisor $tK_{f}+M_{i}-\mu_{i}F+Z_{i}$ is nef for any $t\geq\alpha_{i}$. In particular, $(\alpha_{i}+1)K_{f}-\mu_{i}F\equiv\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i}$ is a nef divisor.
Proof.
For the reader’s convenience we rielaborate here Konno’s proof of [34, Lemma 2.1]. As in the statement we call $Z_{i}=\sum_{j}m_{j}G_{j}$ be the irreducible decomposition of $Z_{i}$, and let $\alpha_{i}:=\max_{j}\{m_{j}\mid F\cdot G_{j}>0\}$; so, $\alpha_{i}$ is the maximal multiplicity of the components of $Z_{i}$ that are not vertical (note that we already know from Remark 3.10 that we can assume that $Z_{i}$ contains no fibres, but it can contain some component of fibres of $f$). We know from (1) of Theorem 2.2 that $K_{f}$ and from Lemma 3.8 and Proposition 3.9 that $M_{i}-\mu_{i}F$ are nef. Let us now consider any irreducible curve $C\subset S$. If $C$ is not a component of $Z_{i}$, then $C\cdot Z_{i}\geq 0$ and so $(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C\geq 0$. Let us now assume that $C=G_{j}$ for some $j$. If $C$ is vertical, $C\cdot F=0$ and so
$(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C=((\alpha_{i}+1)K_{f}-\mu_{i}F)\cdot C=(\alpha_{i}+1)K_{f}\cdot C\geq 0.$
If on the contrary, $C\cdot F>0$, then the restriction of $f$ to $C$ is a finite surjective morphism $\pi=f_{|C}\colon C\to B$ of degree $d=C\cdot F$. We have that
$(K_{f}+C)\cdot C=(K_{S}+C-f^{*}K_{B})\cdot C=\deg\omega_{C}-d(2b-2)=\deg r(\pi)\geq 0,$
by the Riemann-Hurwitz formula, where $r(\pi)$ denotes the ramification divisor of $\pi$ on $B$. Now we can rearrange the divisor $\alpha_{i}K_{f}+Z_{i}$ as follows:
$\alpha_{i}K_{f}+Z_{i}=(\alpha_{i}-m_{j})K_{f}+m_{j}(K_{f}+C)+(Z_{i}-m_{j}C);$
and so we have
$(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C=(\alpha_{i}-m_{j})K_{f}\cdot C+m_{j}(K_{f}+C)\cdot C+(Z_{i}-m_{j}C)\cdot C+(M_{i}-\mu_{i}F)\cdot C,$
and one readily verifies that all the contributions are greater or equal to zero, by making the following observations:
  • (i)
    $\alpha_{i}-m_{j}\geq 0$ by definition of $\alpha_{i}$, and $K_{f}$ is nef.
  • (ii)
    $m_{j}(K_{f}+C)\cdot C=m_{j}r(\pi)\geq 0$.
  • (iii)
    $Z_{i}-m_{j}C$ is an effective divisor which does not contain $C$ in its support, so it has non-negative intersection with $C$.
  • (iv)
    $M_{i}-\mu_{i}F$ is nef.
The proof is thus concluded. ∎
Remark 3.14.
It is important to remark that this idea of Konno introduces a real novelty with respect to Xiao’s method; indeed, the latter method constructs nef divisors on $S$ by pulling back (extremal) nef divisors on $\mathbb{P}_{B}(\mathcal{E}_{i})$, so it obtains divisors of the form $M_{i}-\mu_{i}F\equiv K_{f}-Z_{i}-\mu_{i}F$. In particular it does not give informations on the nefness of divisors of the form $K_{f}-bF$, unless $Z_{i}=0$.
Remark 3.15.
Notice that $\alpha_{i}\leq a_{i}$, and if $\alpha_{i}=a_{i}$, then necessarily $Z_{i}=a_{i}\Gamma$, where $\Gamma$ is a section of $f$, and moreover $a_{i}\Gamma$ is contained in all the $Z_{j}$ with $j\leq i$.
Lemma 3.16.
With the above notations, we have that for any $i=1,\ldots l$,
$K_{f}\cdot Z_{i}\geq\mu_{i}\frac{a_{i}}{\alpha_{i}+1}.$ (3.17)
Proof.
We consider the rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$. As observed above, $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$, with $Z_{i}$ effective, and $M_{i}-\mu_{i}F$ is nef. By Proposition 3.13, as of course $a_{i}\geq\alpha_{i}$, we have that $(\alpha_{i}+1)K_{f}-\mu_{i}F\equiv\alpha_{i}K_{f}+M_{i}+Z_{i}-\mu_{i}F$ is a nef divisor. So, $((\alpha_{i}+1)K_{f}-\mu_{i}F)\cdot Z_{1}\geq 0$, and ultimately we get that $K_{f}\cdot Z_{i}\geq a_{i}\mu_{i}/(\alpha_{i}+1)$. ∎
We now follow [34, Lemma 2.7] and make explicit the main inequality proved in that argument:
Lemma 3.18.
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$. Let $a:=a_{1}$, and $\alpha:=\alpha_{1}$. The following inequality holds:
$K_{f}^{2}\geq\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right)\mu_{1}.$ (3.19)
Proof.
We just consider the first rational map $\phi_{1}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{1})$. By Lemma 3.16 we have that $K_{f}\cdot Z_{1}\geq a\mu_{i}/(\alpha+1)$.
Now we use a reasoning very close to the classical Xiao’s trick to bound the self-intersection of $K_{f}$:
$\begin{split}K_{f}^{2}&=(K_{f}\cdot M_{1}+Z_{1})=(K_{f}\cdot M_{1}+K_{f}\cdot Z_{1})=(K_{f}\cdot M_{1}-\mu_{1}F)+(K_{f}\cdot\mu_{1}F)+(K_{f}\cdot Z_{1})=\\ &=(K_{f}-\mu_{1}F\cdot M_{1}-\mu_{1}F)+(\mu_{1}F\cdot M-\mu_{1}F)+(K_{f}\cdot\mu_{1}F)+(K_{f}\cdot Z_{1}).\end{split}$ (3.20)
Now, we observe that $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is an effective divisor, because $M_{1}-\mu_{1}F$ is effective by Theorem 3.2 (as $\mu_{1}\geq\mu_{l}$), and $Z_{1}$ is effective by its very definition. So, the first term $(K_{f}-\mu_{1}F\cdot M_{1}-\mu_{1}F)$ is greater or equal to 00, and we obtain
$\begin{split}K_{f}^{2}\geq\mu_{1}(F\cdot M_{1})+\mu_{1}(K_{f}\cdot F)+(K_{f}\cdot Z_{1})\geq\\ \geq\mu_{1}\left((2g-2-a+2g-2)+\frac{a}{\alpha+1}\right)=\\ =\mu_{1}\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right),\end{split}$ (3.21)
as claimed. ∎
Remark 3.22.
More generally one can prove that for any $i=1,\ldots,l$
$K_{f}^{2}\geq\left(\frac{(\alpha_{i}+1)(4g-4-a_{i})+a_{i}}{\alpha_{i}+1}\right)\mu_{i}.$ (3.23)
In some cases also these inequalities can be useful, as we will see for instance in the last part of next section, in particular in Theorem 4.30.
4 Some new slope inequalities
We prove in this section several slope inequalities involving the unitary rank $u_{f}$, and some also involving $c_{f}$, and the rank $r_{1}$ of the first subbundle $\mathcal{E}_{1}$.
4.1 First applications
Even avoiding the elaboration made in Lemmas 3.16 and 3.8 one can derive directly from Proposition 3.13 some interesting slope inequalities, which will be very important for the applications of the following sections.
Theorem 4.1.
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$. Let $\alpha:=\alpha_{1}$ be the maximal multiplicity of the non-vertical irreducible components of $Z_{1}$, the following inequality holds:
$K_{f}^{2}\geq\left(\frac{2g-2}{\alpha+1}+2g-2\right)\mu_{1}.$ (4.2)
In particular, the following inequality holds:
$K_{f}^{2}\geq\frac{4g(g-1)}{(2g-1)(g-u_{f})}\chi_{f}.$ (4.3)
Moreover, equality holds in (4.3) if and only if $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F=(2g-2)\Gamma+\mu_{1}F$, where $\Gamma$ is a section of $f$.
Proof.
First observe that the assumption that $f$ is non-locally trivial implies that $\chi_{f}>0$ and so equivalently (Remark 3.5) $u_{f}\leq g-1$.
From Konno’s result (Proposition 3.13) we have that $(\alpha+1)K_{f}-\mu_{1}F$ is a nef divisor on $S$, and by Xiao’s result (Proposition 3.9) $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is a pseudoeffective divisor, so their intersection is greater or equal to 00:
$((\alpha+1)K_{f}-\mu_{1}F\cdot K_{f}-\mu_{1}F)=(\alpha+1)K_{f}^{2}-(2g-2)(\alpha+2)\mu_{1}\geq 0,$
so
$K_{f}^{2}\geq(2g-2)\frac{\alpha+2}{\alpha+1}\mu_{1}\geq\frac{(\alpha+2)(2g-2)}{(\alpha+1)(g-u_{f})}\chi_{f}.$
So, inequality (4.2) is proved.
Now just observe that the function
$t\mapsto\frac{t+2}{t+1}$
is strictly decreasing and $\alpha\leq 2g-2$ so we obtain inequality (4.3).
As for the case where equality holds, we see that we need to have equalities in all the passages of the proof, so in particular $\alpha=a=2g-2$, i.e $d_{1}=0$, so necessarily we have (see Remark 3.7) $r_{1}=1$. Moreover, we need to have $\mu_{1}=\chi_{f}/(g-u_{f})=\mu(\mathcal{A})$, but this only can happen if $\mathcal{A}$ is semistable (see $(ii)$ of Remark 3.6) hence $l=2$. Putting all together, we have that necessarily $g-u_{f}=\mbox{{rank}}\mathcal{A}=\mbox{{rank}}\mathcal{E}_{1}=1$ and $\mu_{1}=\chi_{f}$. Note moreover that, as $\alpha=a=2g-2$ the divisor $Z_{1}$ need to be $(2g-2)\Gamma+Z^{\prime}$, where $\Gamma$ is a section of $f$, and $Z^{\prime}$ is a vertical divisor.
Now we prove that necessarily if equality holds in (4.3) then $Z^{\prime}=0$.
First of all we observe that $Z^{\prime}$ does not contain any fibre in its support (see also Remark 3.10): if this was the case, we would have
$K_{f}\equiv(2g-2)\Gamma+bF+Z^{\prime\prime}+\mu_{1}F=(2g-2)\Gamma+Z^{\prime\prime}+(\mu_{1}+b)F,$
with $b\geq 1$ and $Z^{\prime\prime}\geq 0$ vertical, so $K_{f}-(\mu_{1}+b)F$ would be effective and we would have a sharper inequality than (4.3).
We now prove that there are no vertical components. Suppose on the contrary $Z^{\prime}\not=0$, and let $D$ be an irreducible component of $Z^{\prime}$ (which is not a fibre by the previous argument). We have equality in (4.3)
$K_{f}^{2}=\frac{4g(g-1)}{(2g-1)(g-u_{f})}\chi_{f};$
On the other side $K_{f}^{2}=\frac{4g(g-1)}{2g-1}\mu_{1}+4(g-1)\Gamma\cdot Z^{\prime}$, so we need to have $\Gamma\cdot Z^{\prime}=0$ which implies $K_{f}\cdot Z^{\prime}=Z^{\prime 2}$. But $Z^{\prime 2}\leq 0$ because $Z^{\prime}$ is vertical, and so $K_{f}\cdot Z^{\prime}=Z^{\prime 2}=0$. But the algebraic index theorem [4, Thm. 4.14], as $K_{f}^{2}>0$ (Theorem 2.2) implies that $Z^{\prime 2}<0$assuming $Z^{\prime}\not=0$- which gives the desired contradiction.
So, we eventually get that $K_{f}$ is necessarily numerically equivalent to $(2g-2)\Gamma+\mu_{1}F=(2g-2)\Gamma+\chi_{f}F$.
On the other side, suppose that $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$, with $\Gamma$ a section of $f$. Then $K_{f}^{2}=2(2g-2)\chi_{f}+4(g-1)^{2}\Gamma^{2}$. Now, observe that $K_{f}\cdot\Gamma=-\Gamma^{2}$ by adjunction, and that -by the Algebraic index theorem again- the intersection matrix between $K_{f},\Gamma$ and $F$ has zero determinant:
$0=\det\begin{pmatrix}K_{f}^{2}&K_{f}\cdot\Gamma&K_{f}\cdot F\\ K_{f}\cdot\Gamma&\Gamma^{2}&\Gamma\cdot F\\ K_{f}\cdot F&\Gamma\cdot F&F^{2}\end{pmatrix}=\det\begin{pmatrix}K_{f}^{2}&-\Gamma^{2}&2g-2\\ -\Gamma^{2}&\Gamma^{2}&1\\ 2g-2&1&0\end{pmatrix}=-2g(2g-2)\Gamma^{2}-K_{f}^{2}$
and hence $K_{f}^{2}=-4g(g-1)\Gamma^{2}$. Combining the two equalities we obtain precisely
$K_{f}^{2}=\frac{4g(g-1)}{2g-1}\chi_{f}.$
and we obtain equality in (4.3). ∎
Remark 4.4.
Inequality (4.3) is a a generalization of [34, Lemma 2.7]. For $u_{f}\ll g$ this result is not very interesting, for instance in case $u_{f}=0$ (4.3) gives as a coefficient $(4(g-1))/(2g-1)$, which is strictly smaller than the one of the slope inequality. But for $u_{f}$ large with respect to $g$ both inequalities become interesting. Indeed we will see in Section 6 that the last one implies a Xiao type inequality on $u_{f}$.
We now see how we can use again the Hodge index theorem already mentioned in the proof of Theorem 4.1 in combination with Konno’s method, to obtain a somewhat different inequality as follows. Let $\Sigma$ be the irreducible horizontal component of $Z_{1}$ of maximal multiplicity $\alpha_{=}\alpha_{1}$. If there is more than one such components we simply choose one. Then the restriction of $f$ to $\Sigma$ is a (ramified) covering of $\pi\colon\Sigma\to B$, of degree say $\beta$.
Theorem 4.5.
With the above notations, the following inequality holds:
$K_{f}^{2}\geq 4\frac{(4g-4-a)(g-1)(g-1+\beta)}{4(g-1)(g+1+\beta)-\beta^{2}\alpha}\mu_{1}$ (4.6)
Proof.
First of all observe that $K_{f}\cdot\Sigma+\Sigma^{2}=r\pi\geq 0$. As $K_{f}^{2}>0$ we have that the intersection matrix between $K_{f},\Sigma$ and $F$ has determinant greater or equal to 00:
$0\leq\det\begin{pmatrix}K_{f}^{2}&K_{f}\cdot\Sigma&K_{f}\cdot F\\ K_{f}\cdot\Sigma&\Sigma^{2}&\Sigma\cdot F\\ K_{f}\cdot F&\Sigma\cdot F&F^{2}\end{pmatrix}=\det\begin{pmatrix}K_{f}^{2}&K_{f}\cdot\Sigma&2g-2\\ K_{f}\cdot\Sigma&\Gamma^{2}&\beta\\ 2g-2&\beta&0\end{pmatrix}$
so we obtain
$0\leq 4\beta(g-1)K_{f}\cdot\Sigma-4(g-1)^{2}\Sigma^{2}-\beta^{2}K_{f}^{2}\leq(4\beta(g-1)+4(g-1)^{2})K_{f}\cdot\Sigma-\beta^{2}K_{f}^{2},$
and so
$K_{f}\cdot Z_{1}=K_{f}\cdot(\alpha\Sigma+Z^{\prime})\geq\alpha K_{f}\cdot\Sigma\geq\frac{\alpha\beta^{2}}{4(g-1)(\beta+g-1)}K_{f}^{2}.$
Now, plugging this inequality in the argument of Lemma 3.18, and in particular in (3.21), we obtain:
$K_{f}^{2}\geq(4g-4-a)\mu_{1}+K_{f}\cdot Z_{1}\geq(4g-4-a)\mu_{1}+\frac{\alpha\beta^{2}}{4(g-1)(\beta+g-1)}K_{f}^{2};$
and so we obtain the desired inequality. ∎
Remark 4.7.
Observe that in case $a=\alpha=2g-2$ (and thus $\beta=1$), we obtain (4.3). In general inequality (4.6) seems to be weaker than (4.2), unless one has specific informations on $\beta$ and $\alpha$.
Theorem 4.8.
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Assume that $Z_{1}$ doesn’t have horizontal components, i.e. $a=\alpha=0$ (equivalently the linear system induced by the stalk of $\mathcal{E}_{1}$ on the general fibre is base point free); Then the following slope inequality holds:
$K_{f}^{2}\geq\frac{4(g-1)}{g-u_{f}}\chi_{f}.$ (4.9)
Proof.
Inequality (4.9) can be deduced from Theorem 4.1 setting $\alpha=0$ in (4.3). ∎
Remark 4.10.
Notice however that inequality (4.9) is false in general, as proved by the examples of Lu in [41]. Moreover, the assumption on the base locus of the system induced by $\mathcal{E}_{1}$ is quite restrictive, and not very easy to verify in concrete cases.
Remark 4.11.
The previous theorem is somewhat close to one due to the author with Barja in [7, Remark 3.3]: there we proved that inequality (4.9) holds if the ample summand $\mathcal{A}$ of the Hodge bundle is semistable and its stalk induces a base point free system on the general fibres. Recall that $\mathcal{A}$ is either the whole $f_{*}\omega_{f}$ or the one to last subsheaf of the Harder-Narashiman sequence $\mathcal{E}_{l}$. So, the assumptions of [7, Remark 3.3] can be rephrased as follows: either the Hodge bundle is semistable (in this case by definition $Z_{1}=0$) or it has a Harder-Narashiman sequence of lenght two, $\mu_{2}=0$ and $Z_{1}$ has no horizontal components. So, the above Theorem implies [7, Remark 3.3].
4.2 Some new slope inequalities deriving from Lemma 3.16
We now use Lemma 3.16, and see that in many cases it gives stronger slope inequalities than the known ones.
Theorem 4.12.
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Let us denote as usual $r_{1}=\mbox{{rank}}\mathcal{E}_{1}$.
The following inequality holds:
$K^{2}_{f}\geq\left(\frac{2g+2r_{1}-4}{g-u_{f}}+\frac{2g-2r_{1}}{(2g-2r_{1}+1)(g-u_{f})}\right)\chi_{f}.$ (4.13)
Moreover, suppose that equality is reached. Then:
  • if $r_{1}=1$ then $f$ is a trigonal fibration;
  • if $r_{1}\geq 2$ then $F$ is a hyperelliptic fibration.
Proof.
If $u_{f}>0$, then by Remark 3.5 we have $\mathcal{A}=\mathcal{E}_{l-1}$, and the Harder-Narashiman sequence of $\mathcal{A}$ is just the truncation of the one of $\mathcal{E}$:
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\ldots\subset\mathcal{E}_{l-1}=\mathcal{A}.$
Let us recall inequality (3.8):
$K_{f}^{2}\geq\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right)\mu_{1}.$
Now, observe that $a\geq\alpha$ and that the function
$f(t):=\frac{(t+1)(4g-4-t)+t}{t+1}$
is strictly decreasing.
If $r_{1}=1$, then $a=2g-2$, and we obtain (4.3), which coincides with inequality (4.13).
If $r_{1}>1$, then we can apply Clifford’s theorem ([3, Chap.III Sec 1]) to the subcanonical linear system ${M_{i}}_{|F}$ and obtain $d_{1}\geq 2(r_{1}-1)$, so $a_{1}=2g-2-d_{1}\leq 2g-2r_{1}$, so we can use the inequality as follows:
$\begin{split}K_{f}^{2}&\geq\mu_{1}f(2g-2r_{1})=\frac{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}{2g-2r_{1}+1}\mu_{1}\geq\\ &\geq\frac{(2g+2r_{1}-4)(2g-2r_{1}+1)+2g-2r_{1}}{2g-2r_{1}+1}\frac{\chi_{f}}{g-u_{f}},\end{split}$
as wanted. Let us now suppose that equality holds in (4.13). If $r_{1}=1$, we proved in Theorem 4.1 that necessarily $u_{f}=g-1$, and so $f$ is a trigonal fibration by Proposition 6.4 below. If $r_{2}\geq 2$ then the fibre of $\mathcal{E}_{1}$ induce a linear series on $F$, of degree necessarily $d_{1}=2r_{1}-2$. So, if $r_{1}<g$, again Clifford’s Theorem tells us that $F$ is hyperelliptic. If $r_{1}=g$, then the Hodge bundle is semistable, and in particular $u_{f}=0$. So, the bound becomes the classical slope inequality (2.15), and we can use Theorem 2.14. ∎
Remark 4.14.
Notice that in the worst case possible for (1) i.e. for $r_{1}=1$, we obtain again precisely the bound (4.3) of Theorem 4.1.
We can somehow improve this inequality by using a sharpening of Clifford’s theorem obtained by the author with E. Riva in [48, Theorem 2.13]. Let us recall the result in the form needed here:
Theorem 4.15 (Riva-Stoppino).
Let $F\subseteq\mathbb{P}^{g-1}$ be a canonical non-hyperelliptic curve of genus $g\geq 2$. For any subspace $W\subset H^{0}(K_{F})$ of dimension $\dim W\geq 2$ and codimension $k$, we have:
$\frac{\deg|W|}{\dim|W|}\geq\frac{2g-2-\beta}{g-\beta-1},$
where $\beta:=\min\{k,\mbox{{Cliff}}(C)\}$.
Theorem 4.16.
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Let $m:=min\{c_{f},g-r_{1}\}$. The following inequality holds:
$K^{2}_{f}\geq\left(2g-1+\frac{2g-2-m}{g-m-1}(r_{1}-1)-\frac{g-m-1}{(2g-1)(g-m-1)-(2g-2-m)(r_{1}-1)}\right)\frac{\chi_{f}}{g-u_{f}}.$ (4.17)
Proof.
First two simple remarks. If the fibration is hyperelliptic i.e. $c_{f}=0$, we have $m=0$ and so inequality (4.17) becomes precisely inequality (4.13) of the theorem above. If $r_{1}=1$, then inequality (4.17) again becomes (4.3).
So we have reduced ourselves to the case when $F$ is non-hyperelliptic and $r_{1}\geq 2$. In this case we are in the position to apply Theorem 4.15, and the theorem follows immediately by applying Lemma 3.18 observing that
$\alpha\leq a\leq 2g-2-\frac{2g-2-m}{g-m-1}(r_{1}-1).$
Remark 4.18.
An interesting application of this result is the consequences we can derive in the case that $\mathcal{A}$ is semistable: see the next section.
In case $r_{1}\geq 2$ we can give sharper inequalities as follows:
Theorem 4.19.
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Let us suppose that $r_{1}=\mbox{{rank}}\mathcal{E}_{1}\geq 2$. Consider the special divisor $R_{1}:=(K_{f}-Z_{1})_{|F}$ on the general fibre $F$.
Then one of the following possibilities hold:
  • (i)
    If $h^{1}(R_{1})\geq 2$, then
    $\begin{split}K_{f}^{2}&\geq\left(\frac{2g+2r_{1}+c_{f}-4}{g-u_{f}}+\frac{2g-2r_{1}-c_{f}}{(2g-2r_{1}-c_{f}+1)(g-u_{f})}\right)\chi_{f}\geq\\ &\geq\left(\frac{2g+c_{f}}{g-u_{f}}+\frac{2g-4-c_{f}}{(2g-3-c_{f})(g-u_{f})}\right)\chi_{f}.\\ \end{split}$ (4.20)
  • (ii)
    If $h^{1}(R_{1})=1$, then necessarily $R_{1}\not=0$ and the following inequality holds:
    $K_{f}^{2}\geq\left(\frac{3g+r_{1}-4}{g-u_{f}}+\frac{g-r_{1}}{(g-r_{1}+1)(g-u_{f})}\right)\chi_{f}\geq\left(\frac{3g-1}{g-u_{f}}+\frac{g-2}{(g-1)(g-u_{f})}\right)\chi_{f}.$ (4.21)
Proof.
Observe that $R_{1}$ is a special divisor of degree $d_{1}$ and ${M_{1}}_{|F}\subseteq H^{0}(F,R_{1})$, and so $h^{0}(R_{1})\geq r_{1}\geq 2$ by assumption.
(i) If $h^{1}(R_{1})\geq 2$, then $R_{1}$ contributes to the Clifford index (REF) of $F$ and so $d_{1}\geq 2(h^{0}(R_{1})-1)+\mbox{{Cliff}}(F)\geq 2r_{1}-2+c_{f}$. So we have that $a_{1}\leq 2g-2r_{1}-c_{f}$, and we obtain equation (4.20) by using (3.19).
(ii) If $h^{1}(R_{1})=1$, in theory $R_{1}$ could be 00, but in this case we would have necessarily $d_{1}=0$ and so $r_{1}=1$, contrary to the assumption. So, necessarily $R_{1}\not=0$, and by Riemann-Roch
$d_{1}-g+2=h^{0}(D_{1})\geq r_{1},$
hence $a_{1}\leq g-r_{1}$, and we just use this bound again in (3.19). ∎
4.3 Slope inequalities for Harder-Narashiman sequence of lenght $\leq 2$
Now we prove new slope inequalities for the case where the Harder-Narashiman sequence is short. For the case $\ell=1$, i.e $f_{*}\omega_{f}$ is semistable, recall the following beautiful result:
Theorem 4.22 (Lu-Zuo [39]).
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$. If the Hodge bundle is semistable, then the following inequality holds:
$K_{f}^{2}\geq\frac{5g-6}{g}\chi_{f}.$ (4.23)
Proof.
We give the proof because it is very short: being $f_{*}\omega_{f}$ semistable, so is its symmetric product $\mbox{{Sym}}^{2}f_{*}\omega_{f}$. Consider the relative multiplication morphism $\gamma\colon\mbox{{Sym}}^{2}f_{*}\omega_{f}\to f_{*}\omega_{f}^{\otimes 2}$. By Noether’s theorem, $im\gamma$ has the same rank of $f_{*}\omega_{f}^{\otimes 2}$, so
$\mu(\mbox{{Sym}}^{2}f_{*}\omega_{f})\leq\mu(im\gamma)\leq\mu(f_{*}\omega_{f}^{\otimes 2}).$
But $\deg f_{*}\omega_{f}^{\otimes 2}=K_{f}^{2}+\chi_{f}$ (by Grothendieck-Riemann-Roch), and $\mu(\mbox{{Sym}}^{2}f_{*}\omega_{f})=2\mu(f_{*}\omega_{f})=2\chi_{f}/g$. So, the inequality above becomes
$2\frac{\chi_{f}}{g}\leq\frac{K_{f}^{2}+\chi_{f}}{3g-3},$
which gives the statement. ∎
Remark 4.24.
The result above gives a high bound for the case in which the Harder-Narasimhan sequence is of lenght $1$. It is not known wether this bound is sharp or not.
We are now going to give sharp bounds for the cases where the Harder-Narasimhan sequence of $\mathcal{E}=f_{*}\omega_{f}$ is of lenght $2$.
First we treat the case when $\mu_{l}=0$ (i.e. $u_{f}>0$).
Theorem 4.25.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$, such that $\mathcal{A}$ is semistable and $\mu_{l}=0$ (so $u_{f}>1$). Then the following inequality holds:
$K_{f}^{2}\geq\left(\frac{4(g-1)-2u_{f}}{(g-u_{f})}+\frac{2u_{f}}{(2u_{f}+1)(g-u_{f})}\right)\chi_{f}.$ (4.26)
Moreover, assume that equality holds in (4.26); then:
    if $u_{f}\leq g-2$, then $F$ is hyperelliptic;
    if $u_{f}=g-1$, then $F$ is trigonal.
Assume that $u_{f}\leq c_{f}$. Then Then the following inequality holds:
$K_{f}^{2}\geq\left(\frac{4(g-1)-u_{f}}{(g-u_{f})}+\frac{u_{f}}{(u_{f}+1)(g-u_{f})}\right)\chi_{f}.$ (4.27)
Assume that $u_{f}\geq c_{f}$. Then the following inequality holds:
$K_{f}^{2}\geq\left(2g-1+\frac{2g-2-c_{f}}{g-c_{f}-1}(g-u_{f}-1)-\frac{g-c_{f}-1}{g(2u_{f}-2c_{f}+1)+u_{f}(g-c_{f}-2)-2}\right)\frac{\chi_{f}}{(g-u_{f})}.$ (4.28)
Proof.
By the assumptions the Harder-Narasimhan sequence of the Hodge bundle $\mathcal{E}$ has lenght 2 and is of this form:
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}=\mathcal{A}\subset\mathcal{E}_{2}=\mathcal{E}.$
Moreover, $\mu_{1}=\chi_{f}/(g-u_{f})$ and $\mu_{2}=0$. Observe that under our assumptions $r_{1}=g-u_{f}$.
We use Theorem 4.12, observing that under our assumptions $r_{1}=g-u_{f}$, and (4.13) becomes
$K^{2}_{f}\geq\left(\frac{4g-4-2u_{f}}{g-u_{f}}+\frac{2u_{f}}{(2u_{f}+1)(g-u_{f})}\right)\chi_{f},$
as wanted.
The case where equality holds follows again from Theorem 4.12. As for the other two inequalities, they follow directly from the result of Theorem 4.16. ∎
Remark 4.29.
Observe that using Xiao’s method with the assumptions of Theorem 4.25 one obtains in this case (as explained for instance in [7, Remark 3.3]) the worse inequality:
$K_{f}^{2}\geq\left(\frac{4(g-1)-2u_{f}}{(g-u_{f})}\right)\chi_{f}.$
This very same equality is reproved in the article [17] in the proof of Lemma 3.3.2.
Although we will not present any consequence of this result, let us complete the picture by giving an inequality for the case when the Harder-Narashiman sequence has lenght 2 and $u_{f}=0$.
Proposition 4.30.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$, such that $l=2$, and $\mu_{l}>0$ (so in particular $u_{f}=0$). Then the following inequality holds:
$K_{f}^{2}\geq\frac{(4(g-1)[(2g-2r_{1}+1)(g+r_{1}-2)+g-r_{1}}{2g^{3}+(2r_{1}-2)g^{2}+(-6{r_{1}}^{2}+9r_{1}-2)g+2{r_{1}}^{2}-3{r_{1}}^{2}}\chi_{f}.$ (4.31)
Proof.
We have that $r_{2}=g$, and (3.1) $\chi_{f}=\mu_{1}r_{1}+\mu_{2}(r_{2}-r_{1})$. By Lemma 3.16 and Remark 3.22 we have
$K_{f}^{2}\geq\left(\frac{(\alpha_{i}+1)(4g-4-a_{i})+a_{i}}{\alpha_{i}+1}\right)\mu_{i},$
for $i=1,2$.
For $i=1$ we have that $\alpha_{1}\leq a_{1}\leq 2g-2r_{1}$, so
$K_{f}^{2}\geq\frac{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}{2g-2r_{1}+1}\mu_{1}.$
For $i=2$ we have $a_{2}=\alpha_{2}=0$, so we obtain
$K_{f}^{2}\geq 4(g-1)\mu_{2}$
So, putting these two inequalities together, we obtain
$\chi_{f}=\mu_{1}r_{1}+\mu_{2}(g-r_{1})\leq\left(\frac{2g-2r_{1}+1}{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}r_{1}+\frac{g-r_{1}}{4(g-1)}\right)K_{f}^{2}$
and a straightforward computation gives the desired inequality. ∎
Remark 4.32.
The coefficient of $\chi_{f}$ is an increasing function of $r_{1}$, in the worst case $r_{1}=1$ we have
$K_{f}^{2}\geq\frac{8g(g-1)^{2}}{2g^{3}+g-1}\chi_{f},$
which is worse than the slope inequality. But for increasing values of $r_{1}$ the function gets better. For instance, in the best case $r_{1}=g-1$ we have
$K_{f}^{2}\geq\left(\frac{3(g-1)}{2}+\frac{1}{4(g-1)}\right)\chi_{f},$
which of course is extremely high.
5 A discussion on the extremal cases
We now list some results, due to Serrano, Pirola, Zucconi and Beorchia, about non locally trivial fibred surfaces with maximal unitary rank $u_{f}=g-1$
First of all, in [47, Theorem 3.6 and Remark 3.7] Pirola proved that in this case up to a finite base change $g-1=u_{f}=q_{f}$ (see Remark 2.23 to see that this is not obvious).
Proposition 5.1.
Let $f\colon S\to B$ a relatively minimal non trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the general fibre $F$ is trigonal;
Proof.
By Pirola’s result we can assume that the relative irregularity is maximal. Cai proved in [15] that for non locally trivial hyperelliptic fibrations Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds, so for $g\geq 3$ no hyperelliptic fibration can have $q_{f}=g-1$. In case $g=2$ of course $F$ is hyperelliptic. The case $g=3$ has been investigated in [44]. On the other hand, by Barja-Naranjo-Gonzalez [6] (later generalized to $u_{f}$ in [30]), we have that $q_{f}\leq g-c_{f}$ where $c_{f}=\mbox{{Cliff}}(F)$ is the Clifford index of the general fibre. Moreover Favale-Naranjo-Pirola [24] proved that if the general fibre is a degree $d$ plane curve (thus of gonality $d-1$ and Clifford index $d-3$ [3]), then $q_{f}\leq g-c_{f}-1$. So, under the assumption of the proposition, $\mbox{{Cliff}}(F)\leq 1$ and $F$ is not a plane quintic, nor it is hyperelliptic (i.e. with $\mbox{{Cliff}}(F)>0$). Hence we conclude that $F$ necessarily is trigonal. Observe that for this last part we could have used directly the unitary rank using the results of the author with Gonzàles-Alonso and Torelli in [30]. ∎
Theorem 5.2.
Let $f\colon S\to B$ a relatively minimal non trivial fibred surface of genus $g\geq 4$ such that $u_{f}=g-1$. Then the following hold:
  1. (1)
    the fibration is non-isotrivial;
  2. (2)
    the smooth fibres $F_{t}$ are a covering of a non-isotrivial family of elliptic curves $E_{t}$.
  3. (3)
    the fibration is not Kodaira , i.e. $K_{f}^{2}<12\chi_{f}$.
Proof.
(1) This is due to Serrano: indeed, in [49] he proves that Xiao’s bound (2.18) $q_{f}\leq(g+1)/2$ holds for isotrivial fibrations. So we can exclude the isotrivial fibrations unless $g\leq 3$. In in [44] are indeed constructed also isotrivial fibrations of genus $2$ and $3$ with irregularity respectively $1$ and $2$.
(2) This proof is due to Pietro Pirola in a private communication. Let us consider a smooth fibre of $f$, $F_{t}$, for $t\in B^{0}$ (the smooth locus of $f$). We have the following diagram
[…]
Now, the morphism ${j_{t}}_{*}$ goes in a family of translates of a fixed variety $A\subset\mbox{{Alb}}(S)$ of dimension $q_{f}=u_{f}=g-1$. Let us call $g_{t}\colon J(F_{t})\to A$ this morphism, fixing $A$. Then if we pass to the dual varieties and morphisms we have $(g_{t})^{*}\colon\widehat{A}\hookrightarrow\mbox{{Pic}}^{0}(F_{t})$. Let us consider the quotient abelian variety, which is an elliptic curve, and denote it $E_{t}$. We have the morphisms
$F_{t}\hookrightarrow\mbox{{Pic}}^{0}(F_{t})\longrightarrow E_{t},$
which is necessarily surjective because the image of $F_{t}$ generates $\mbox{{Pic}}^{0}(F_{t})$ as an abelian variety. So, this is the desired morphism of the general fibre to a family of elliptic curves.
Now, the family $E_{t}$ has varying moduli: if this was not the case than all $\mbox{{Pic}}^{0}(F_{t})$ would be a fixed abelian variety and so the entire family would be trivial up to a base change, hence isotrivial, but this is not the case for (2).
(3) Also this proof is due to Pietro Pirola in a private communication. By performing a base change $B^{\prime}\to B$ if necessary, we have a global non-isotrivial family of genus $1$ curves $W\stackrel{{\scriptstyle\alpha}}{{\to}}B^{\prime}$ such that, calling $f^{\prime}\colon S^{\prime}\to B^{\prime}$ the fibration obtained by base change, there exists a covering $\beta\colon S^{\prime}\to W$ such that $f^{\prime}=\alpha\circ\beta$.
${S^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{\beta}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B}$
Now the fibration of genus one contains at least one fibre with non compact Jacobian. Thus $f$ contains at least a fibre with the same property, and this implies that it is not isotrivial.
The last conclusion could be derived also from the fact that a non-isotrivial family of curves of genus $1$ contains at least a vertical rational curve: see [2]. So necessarily $\alpha$ contains fibres whose support has more than one component. This implies that also $f^{\prime}$ necessarily has the same feature, and so also $f$ necessarily contains singular fibres. ∎
6 An upper bound on the unitary rank
We now extend Konno’s bound [34, Lemma 2.7] on $q_{f}$ to $u_{f}$, and improve it obtaining a strict inequality.
Theorem 6.1.
Let $f\colon S\to B$ be a relatively minimal non-isotrivial fibration of genus $g\geq 2$. Then
$u_{f}<g\frac{5g-2}{6g-3}$ (6.2)
Proof.
The non-sharp inequality is immediate by by combining inequalities (4.3) and (2.11). Suppose that the equality equality holds in (6.2), then and equality in (4.3) holds, so $f$ is necessarily of maximal unitary rank $u_{f}=g-1$ and equality in (2.11) holds, so $f$ is necessarily Kodaira. We have seen that this is impossible in Theorem 5.2, thus the inequality is always strict. ∎
Theorem 6.3.
Let $f\colon S\to B$ be a relatively minimal non-isotrivial fibration of genus $g\geq 2$. Then Let us suppose that $u_{f}$ is maximal, i.e. $u_{f}=g-1$. Then $g\leq 6$
Proof.
Inequality (4.3) in the first case becomes
$K_{f}^{2}\geq\frac{4g(g-1)}{2g-1}\chi_{f}.$
If we combine it with (2.11), we obtain $4g(g-1)\leq 3(2g-1)$, so $g^{2}-7g+3\leq 0$, which gives that $g\leq 6$. ∎
We can also give a bound in the case $u_{f}=g-2$. By using equation (6.2) in this case we obtain the equation
$g^{2}-13g+7\leq 0,$
which implies $g\leq 12$, but we now see that we can rule out case $g=12$ as follows.
Proposition 6.4.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$.
Suppose that $u_{f}=g-2$. Then the genus of the fibration is $g\leq 11$.
Proof.
Inequality (4.3) becomes
$K_{f}^{2}\geq\frac{2g(g-1)}{2g-1}\chi_{f},$
and so, combining again with (2.11) we get $g\leq 12$. But for $g=12$ we obtain that equality is reached in (4.3), so $u_{f}$ should be $g-1$, a contradiction. ∎
Remark 6.5.
One could also use the results of Section 5 to exclude the case $g=12$.
I the last section 8 we give many conditions on the case when $g=6$ and $u_{f}=g-1=5$.
7 An application to the Coleman-Oort conjecture
We will use the setting and the notations of the papers [40] and [17]. Let us consider $\mathcal{M}_{g}=\mathcal{M}_{g,\ell}$ the moduli space of of smooth projective curves of genus $g\geq 2$ with a full level $\ell-$structure. and $\mathcal{A}_{g}=\mathcal{A}_{g,\ell}$ the moduli space of $g$-dimensional principally polarized abelian varieties with a full level $\ell$-structure. The level $\ell\geq 3$ assures that these are moduli schemes representing the corresponding moduli functor. Let
$j^{\circ}\colon\mathcal{M}_{g}\longrightarrow\mathcal{A}_{g}$
be the Torelli morphism. The Torelli locus $\mathcal{T}_{g}$ is the closure in $\mathcal{A}_{g}$ of the image of $j^{\circ}$, which we call $\mathcal{T}_{g}^{\circ}$.
A closed positive-dimensional subvariety $Z\subseteq\mathcal{A}_{g}$ is said to be contained generically in the Torelli locus if $Z\subset\mathcal{T}_{g}$ and $Z\cap\mathcal{T}_{g}^{\circ}\not=\emptyset$.
The moduli space $\mathcal{A}_{g}$ is isomorphic to a Shimura variety. In $\mathcal{A}_{g}$ there are Shimura subvarieties. The famous Coleman-Oort conjecture predicts that for large $g$ there do not exist any Shimura subvariety generically contained in the Torelli locus.
For $g\leq 4$ such subvarieties do exist: see [27]. Many authors have proved results in the direction of this conjecture, let us recall at least [28], [18], [29], [40].
Recall that, given $C\subset\mathcal{A}_{g}$ a smooth closed curve, the canonical Higgs bundle $\mathcal{E}_{C}$ on $C$ is the Hodge bundle given by the universal family of abelian varieties restricted over $C$. The bundle $\mathcal{E}_{C}$ decomposes as $\mathcal{E}_{C}=\mathcal{F}_{C}\oplus\mathcal{U}_{C}$ where $\mathcal{U}_{C}$ is the maximal unitary Higgs subbundle corresponding to the maximal sub-representation on which $\pi_{1}(C)$ acts through a compact unitary group. The bundles $\mathcal{E}_{C}$, $\mathcal{F}_{C}$ and $\mathcal{U}_{C}$ have an Hodge decomposition into a $(-1,0)$ and $(0,-1)$ part, which we call $\mathcal{U}_{C}^{-1,0}$ and $\mathcal{U}_{C}^{0,-1}$.
Here, using the results of the previous sections, we can give an improvement of a result of Chen, Lu and Zuo in [17, Thm 1.1.2], giving a condition on the rank of the $(-1,0)$ part of the maximal unitary Higgs subbundle $\mathcal{U}_{C}$ of a curve $C\subset\mathcal{A}_{g}$ that implies that it is not generically contained in the Torelli locus.
Theorem 7.1.
Let $C\subset\mathcal{A}_{g}$ be a curve with Higgs bundle decomposition $\mathcal{E}_{C}=\mathcal{A}_{C}\oplus\mathcal{U}_{C}$, where $\mathcal{U}_{C}$ is the maximal unitary Higgs subbundle.
  • (i)
    If
    $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}\geq g\frac{5g-2}{6g-3},$ (7.2)
    then $C$ is not contained generically in the Torelli locus.
  • (ii)
    If $C$ is a Shimura curve such that
    $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}\geq\frac{4g-1+\sqrt{16g^{2}-36g+21}}{10},$ (7.3)
    then $C$ is not contained generically in the Torelli locus.
Proof.
The proof is completely analogous to the one of [17, Thm 1.12]. The key point is that if $C$ is contained generically in the Torelli locus, then we can consider the following diagram
[…]
where $B$ is the normalization of the pullback $(j^{\circ})^{-1}C$ in $\mathcal{M}_{g}$. The morphism $B\to\mathcal{M}_{g}$ gives a holomorphic fibration $f\colon S\to B$ induced by the fine moduli structure. We can perform the semistable reduction to a compactification obtaining a non-isotrivial fibred surface $\overline{f}\colon\overline{S}\to\overline{B}$, and a morphism $\overline{B}\to\overline{\mathcal{M}}_{g}$ that extends $B\to\mathcal{M}_{g}$.
Now, the Hodge bundle $\overline{f}_{*}\omega_{\overline{f}}$ is the $(-1,0)-$part of the Higgs bundle on $\overline{B}$ associated to the relative semistable Jacobian family (see [38, Sec. 3]). The pullback via $j_{\overline{B}}$ of the $(-1,0)$ part of the decomposition $\mathcal{E}_{\overline{B}}=\mathcal{F}_{\overline{B}}\oplus\mathcal{U}_{\overline{B}}$ is precisely the second Fujita decomposition (2.7) of ${\overline{f}}_{*}\omega_{\overline{f}}$, so
$\mbox{{rank}}\,{\mathcal{U}_{\overline{B}}}^{-1,0}=u_{\overline{f}}.$
Now the first statement $(i)$ directly follows from Theorem 6.1. For the second statement $(ii)$ one needs to use a result due to Viehweg and Zuo ([50], see also [17, Thm. 1.2.1]), that tells us in particular that for the case when $C$ is Shimura, then $\mathcal{F}_{\overline{B}}^{-1,0}$ is semistable. So, in this case we are in the assumptions of Theorem 4.25, and so we have inequality (4.26), that combined with (2.11) gives:
$\frac{(4(g-1)-2u_{\overline{f}})(2u_{\overline{f}}+1)+2u_{\overline{f}}}{(2u_{\overline{f}}+1)(g-u_{\overline{f}})}\leq 12\iff$
$(2u_{\overline{f}}+1)(g-1)-u_{\overline{f}}^{2}\leq 3(g-u_{\overline{f}})(2u_{\overline{f}}+1)\iff$
$5u_{\overline{f}}^{2}-(4g-1)u_{\overline{f}}-(2g+1)\leq 0$
Now, the discriminant of this quadratic form is $\Delta=16g^{2}-36g+21>0$, and so we need to have
$u_{\overline{f}}\leq\frac{4g-1+\sqrt{16g^{2}-36g+21}}{10}.$ (7.4)
If equality holds, then in particular $f$ is Kodaira, and equality holds in (4.26), so by Theorem 4.25 we have that:
  • either $u_{\overline{f}}=g-1$, but in this case we already proved in Theorem 5.2 that the fibration is not Kodaira;
  • or $u_{\overline{f}}\leq g-2$ and the fibration is hyperelliptic. In this case we know (see (U4) of [30]) that $u_{\overline{f}}\leq(g+1)/2$, so we get that this is impossible for $g\leq 4$. But also for $g=2,3$ the equality in (7.4) cannot hold, because the discriminant is not a square number.
Remark 7.5.
The bounds found by Chen, Lu and Zuo in [17, Thm 1.12] are respectively:
  • (i)
    if $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}>\frac{5g+1}{6},$ then $C$ is not generically contained in the Torelli locus;
  • (ii)
    if $\mbox{{rank}}\,\mathcal{U}_{C}^{-1,0}\geq\frac{4g+2}{5}$ and $C$ is Shimura, then it is not generically contained in the Torelli locus.
So, as $g\frac{5g-2}{6g-3}<\frac{5g+1}{6}$ and $\frac{4g-1+\sqrt{16g^{2}-36g+21}}{10}<\frac{4g-3}{5}<\frac{4g+2}{5}$, we see that we are giving an improvement to their result. It has also to be noted that in [40] Lu and Zuo prove that there do not exist two specific types of Shimura curves generically contained in $\mathcal{T}_{g}$ for $g\geq 11$.
8 An analysis of the cases $g=6$, $u_{f}=5$, and $g=5$, $u_{f}=4$
We now can be more explicit for what concerns the extremal cases. Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g=6$ such that $u_{f}=5$. First recall that Beorchia Pirola and Zucconi in [13] have proved that such $f$ is a trigonal fibration with special Maroni invariant.
Proposition 8.1.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g=6$ such that $u_{f}=5$. Then $S$ has positive index $\tau(S)>0$, and, calling as usual $\alpha$ the maximal multiplicity of the non vertical components of $Z_{1}$, we have $\alpha\geq 5$.
Proof.
We first prove the statements about the numerical equivalence of $K_{f}$. The assumption tells us that the Harder-Narasimhan sequence of the Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$ is
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}=\mathcal{A}\subset\mathcal{E}_{2}=\mathcal{E},$
and $r_{1}=\mbox{{rank}}\mathcal{E}_{1}=1$. This implies that the map $\phi_{1}\colon S\to\mathbb{P}_{B}(\mathcal{E}_{1})\cong B$ is indeed a morphism and coincides with the fibration $f$. The tautological divisor on $\mathbb{P}_{B}(\mathcal{E}_{1})$ is precisely $\mathcal{E}_{1}=\mathcal{A}$ itself, and $M_{1}=f^{*}\mathcal{A}$. So from Lemma 3.8, we have that $K_{f}\equiv f^{*}\mathcal{A}+Z_{1}$.
Using (4.2), we obtain
$K_{f}^{2}\geq\left(2g-2+\frac{2g-2}{\alpha+1}\right)\mu_{1}\geq\left(10+\frac{10}{\alpha+1}\right)\chi_{f}.$
This implies that the index of $S$ is necessarily positive.
Moreover, if $\alpha\leq 3$, the inequality above would contradict (2.11). If $\alpha=4$, we would have $K_{f}^{2}\geq 12\chi_{f}$, so necessarily equality holds and $f$ is a Kodaira fibration, but this contradicts point (4) of Theorem 5.2
Remark 8.2.
So, we have obtained that $Z_{1}$ has an horizontal component $\Gamma$ of multiplicity greater or equal to $5$. Clearly if $\alpha\geq 6$ the necessarily this component is a section of $f$. In the case $\alpha=5$ we could have that $Z_{1}=5\Sigma$, where $\Sigma$ is a $2:1$ covering of $B$. Unfortunately we have not been able to exclude this case, see also Remark 8.9.
We now see that we can be much more restrictive on the possibilities for the divisor $Z_{1}$. First of all we see that in the extremal case $u_{f}=g-1$, we can improve quite strongly Konno’s bound on $K_{f}\cdot Z_{1}$ in Lemma 3.16 via a direct computation as follows:
Proposition 8.3.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g\geq 2$ such that $u_{f}=g-1$. Let $Z_{1}=\sum^{k}_{j=1}m_{j}G_{j}+Z^{\prime}$, where $G_{j}$ are the irreducible components of $Z_{1}$ which are not vertical with respect to $f$, $m_{j}$ are their respective multiplicities and $z^{\prime}$ is the set of vertical components. Let $b_{j}:=G_{j}\cdot F$. Then we have the following inequalities:
$K_{f}\cdot Z_{1}\geq\left(\sum_{j=1}^{k}\frac{m_{j}}{m_{j}+1}b_{j}\right)\mu_{1}.$ (8.4)
$K_{f}^{2}\geq\left(2(g-1)+\sum_{j=1}^{k}\frac{m_{j}}{m_{j}+1}b_{j}\right)\mu_{1}.$ (8.5)
Proof.
Recall that under our assumptions $f_{*}\omega_{f}=\mathcal{A}\oplus{\mathcal{O}_{B}}^{\oplus(g-1)}$, and $K_{f}\equiv Z_{1}+f^{*}\mathcal{A}\equiv Z_{1}+\mu_{1}F$, and $\mu_{1}=\chi_{f}$.
First of all let us observe that $\deg f_{*}\mathcal{O}_{S}(Z_{1})=-(g-1)\mu_{1}$. This can be proved for instance by observing that
$f_{*}\omega_{f}=f_{*}(\mathcal{O}_{S}(Z_{1}))\otimes f^{*}\mathcal{A})\cong f_{*}\mathcal{O}_{S}(Z_{1})\otimes\mathcal{A}.$
So the formula for the degree of vector bundles gives:
$\deg f_{*}\mathcal{O}_{S}(Z_{1})=-g\deg\mathcal{A}+\deg f_{*}\omega_{f}=-g\chi_{f}+\chi_{f}=-(g-1)\chi_{f}=-(g-1)\mu_{1},$
as wanted.
Now observe that by the relative Serre duality we have
$R^{1}f_{*}(\mathcal{O}_{S}(Z_{1}))\cong(R^{1}f_{*}\omega_{f})\otimes\mathcal{A}^{-1}\cong\mathcal{O}_{S}\otimes\mathcal{A}^{-1}\cong\mathcal{A}^{-1},$
so $\deg R^{1}f_{*}(\mathcal{O}_{S}(Z_{1}))=-\chi_{f}=-\mu_{1}$. Now let us use Grothendieck-Riemann-Roch for $Z_{1}$:
$\deg f_{*}\mathcal{O}_{S}(Z_{1})=\frac{{Z_{1}}^{2}}{2}-\frac{K_{f}\cdot Z_{1}}{2}+\chi_{f}+\deg R^{1}f_{*}(\mathcal{O}_{S}(Z_{1}))=\frac{{Z_{1}}^{2}}{2}-\frac{K_{f}\cdot Z_{1}}{2}.$
Summing up we have
${Z_{1}}^{2}=K_{f}\cdot Z_{1}-2(g-1)\mu_{1}.$ (8.6)
Now we just compute $K_{f}^{2}$ using its numerical equivalence and the equations (8.6):
$K_{f}^{2}={Z_{1}}^{2}+4(g-1)\mu_{1}=K_{f}\cdot Z_{1}+2(g-1)\mu_{1}.$ (8.7)
Next, we want to estimate the number $K_{f}\cdot G_{i}$ for ani $i=1,\ldots,k$. By making explicit the terms, we have
$K_{f}\cdot G_{i}=\left(\sum^{k}_{j=1}m_{j}G_{j}+Z^{\prime}+\mu_{1}F\right)=\sum^{k}_{j=1}m_{j}(G_{j}\cdot G_{i})+(Z^{\prime}\cdot G_{i})+\mu_{1}(F\cdot G_{i})\geq m_{i}G_{i}^{2}+b_{i}\mu_{1}.$
Now, observe that $(K_{f}+G_{i}\cdot G_{i})=\deg r(\pi)$, where $\pi\colon G_{i}\to B$ is the morphism induced by restriction of $f$, and $r(\pi)$ is its ramification divisor. So, $K_{f}\cdot G_{i}\geq-G_{i}^{2}$. Combining this inequality with the above formula, we obtain
$(m_{i}+1)K_{f}\cdot G_{i}\geq b_{i}\mu_{1}.$
Now we just compute $K_{f}^{2}$ using (8.6):
$K_{f}^{2}\geq 2(g-1)\mu_{1}+K_{f}\cdot Z_{1}=2(g-1)\mu_{1}+(K_{f}\cdot\sum^{k}_{j=1}m_{j}G_{j}+Z^{\prime})\geq 2(g-1)\mu_{1}+\left(\sum^{k}_{j=1}\frac{m_{j}b_{j}}{m_{j}+1}\right)\mu_{1},$
as wanted. ∎
Corollary 8.8.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g=6$ such that $u_{f}=5$. The possibilities for the numerical class of $K_{f}$ are the following.
  • (1)
    $\alpha=10$ and $K_{f}\equiv 10\Gamma+\mu_{1}F$ where $\Gamma$ is a section of $f$.
  • (2)
    $\alpha=9$ and $K_{f}\equiv 9\Gamma+\Gamma^{\prime}+\mu_{1}F$ where $\Gamma$ and $\Gamma^{\prime}$ are distinct sections of $f$.
  • (3)
    $\alpha=8$ and
    • either $K_{f}\equiv 8\Gamma+2\Gamma^{\prime}+\mu_{1}F$,
    • or $K_{f}\equiv 8\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\mu_{1}F$, where $\Gamma,\Gamma^{\prime}$ and $\Gamma^{\prime\prime}$ are distinct sections of $f$.
    • or $K_{f}\equiv 8\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection of $f$.
  • (4)
    $\alpha=7$ and $K_{f}\equiv 7\Gamma+3\Gamma+\mu_{1}F$ where $\Gamma$ is a section of $f$,
  • (5)
    $\alpha=6$ and $K_{f}\equiv 6\Gamma+4\Gamma^{\prime}+\mu_{1}F$ where $\Gamma,\Gamma^{\prime}$ are distinct sections of $f$.
  • (6)
    $\alpha=5$ and
    • either $K_{f}\equiv 5\Gamma+5\Gamma^{\prime}+\mu_{1}F$ where $\Gamma,\Gamma^{\prime}$ are distinct sections of $f$,
    • or $K_{f}\equiv 5\Sigma$ where $\Sigma$ is a bisection of $f$.
Proof.
The possible cases according to the values of $\alpha$ are the following. In the list below $\Gamma,\ldots\Gamma^{\prime\prime\prime\prime}$ are distinct sections of $f$, $\Sigma$ is a multisection, i.e. a degree $d>1$ covering of the base $B$. The degree is specified case by case.
  • $(1)^{\prime}$
    $\alpha=10$ and $K_{f}\equiv 10\Gamma+\mu_{1}F$,
  • $(2)^{\prime}$
    $\alpha=9$ and $K_{f}\equiv 9\Gamma+\Gamma^{\prime}+\mu_{1}F$,
  • $(3)^{\prime}$
    $\alpha=8$ and
    • $(3a)^{\prime}$
      $K_{f}\equiv 8\Gamma+2\Gamma^{\prime}+\mu_{1}F$,
    • $(3b)^{\prime}$
      $K_{f}\equiv 8\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\mu_{1}F$,
    • $(3c)^{\prime}$
      $K_{f}\equiv 8\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection of $f$.
  • $(4)^{\prime}$
    $\alpha=7$ and
    • $(4a)^{\prime}$
      $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\mu_{1}F$,
    • $(4b)^{\prime}$
      $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+\mu_{1}F$,
    • $(4c)^{\prime}$
      $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection of $f$,
    • $(4d)^{\prime}$
      $K_{f}\equiv 7\Gamma+3\Gamma^{\prime}+\mu_{1}F$,
    • $(4e)^{\prime}$
      $K_{f}\equiv 7\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a trisection of $f$
  • $(5)^{\prime}$
    $\alpha=6$ and
    • $(5a)^{\prime}$
      $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime}+\mu_{1}F$,
    • $(5b)^{\prime}$
      $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+2\Gamma^{\prime\prime\prime}+\mu_{1}F$,
    • $(5c)^{\prime}$
      $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection of $f$,
    • $(5d)^{\prime}$
      $K_{f}\equiv 6\Gamma+2\Gamma^{\prime}+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection,
    • $(5e)^{\prime}$
      $K_{f}\equiv 6\Gamma+\Sigma+\Sigma^{\prime}+\mu_{1}F$ where $\Sigma,\Sigma^{\prime}$ are distinct bisections,
    • $(5f)^{\prime}$
      $K_{f}\equiv 6\Gamma+3\Gamma+\Gamma^{\prime}+\mu_{1}F$,
    • $(5g)^{\prime}$
      $K_{f}\equiv 6\Gamma+4\Gamma^{\prime}+\mu_{1}F$,
    • $(5h)^{\prime}$
      $K_{f}\equiv 6\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a $4$-section,
  • $(6)^{\prime}$
    $\alpha=5$ and
    • $(6a)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime\prime}+\mu_{1}F$,
    • $(6b)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+2\Gamma^{\prime\prime\prime\prime}+\mu_{1}F$,
    • $(6c)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection,
    • $(6d)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+2\Gamma^{\prime\prime\prime}+\mu_{1}F$,
    • $(6e)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection,
    • $(6f)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma+\Sigma+\Sigma^{\prime}+\mu_{1}F$ where $\Sigma,\Sigma^{\prime}$ are distinct bisections.
    • $(6g)^{\prime}$
      $K_{f}\equiv 5\Gamma+2\Gamma^{\prime}+3\Gamma^{\prime\prime}+\mu_{1}F$,
    • $(6h)^{\prime}$
      $K_{f}\equiv 5\Gamma+2\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a trisection,
    • $(6i)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Sigma+3\Gamma^{\prime}+\mu_{1}F$, where $\Sigma$ is a bisection,
    • $(6j)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+4\Gamma^{\prime\prime}+\mu_{1}F$,
    • $(6k)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a $4$-section,
    • $(6l)^{\prime}$
      $K_{f}\equiv 5\Gamma+5\Gamma^{\prime}+\mu_{1}F$,
    • $(6m)^{\prime}$
      $K_{f}\equiv 5\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ is a $5$-section of $f$.
    • $(6n)^{\prime}$
      $K_{f}\equiv 5\Sigma+\mu_{1}F$ where $\Sigma$ is a bisection of $f$.
So, the possibilities are much more than the ones stated in the theorem. We are going to rule out the great part of these possibilities by using Proposition 6.4.
Let us now consider all the cases we want to exclude:
  1. $(4a)^{\prime}$
    Suppose we have $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\mu_{1}F$.
    So, using inequality 8.5 we obtain
    $K_{f}^{2}\geq\left(10+\frac{7}{8}+\frac{3}{2}\right)\mu_{1}=\left(10+\frac{19}{8}\right)\mu_{1}>12\mu_{1},$
    which contradicts (2.11).
  2. $(4b)^{\prime}$
    Suppose $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+\mu_{1}F$. Using inequality (8.5) we obtain
    $K_{f}^{2}\geq 10\mu_{1}+K_{f}\cdot Z_{1}\geq\left(10+\frac{7}{8}+\frac{1}{2}+1\right)\mu_{1}>12\mu_{1}.$
  3. $(4c)^{\prime}$
    Suppose $K_{f}\equiv 7\Gamma+\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection of $f$. Using inequality (8.5) we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{7}{8}+\frac{1}{2}+1\right)=\left(11+\frac{13}{8}\right)\mu_{1}>12\mu_{1}.$
  4. $(4e)^{\prime}$
    Suppose that $K_{f}\equiv 7\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a trisection of $f$. Using inequality (8.5) we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{7}{8}+\frac{3}{2}\right)\mu_{1}>12\mu_{1}.$
  5. $(5a)^{\prime}$
    Suppose $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime}+\mu_{1}F$. Using inequality (8.5) we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}\right)\mu_{1}>12\mu_{1}.$
  6. $(5b)^{\prime}$
    Suppose $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+2\Gamma^{\prime\prime\prime}+\mu_{1}F$. Using inequality (8.5) we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+\frac{1}{2}+\frac{1}{2}+\frac{1}{3}\right)\mu_{1}>12\mu_{1}.$
  7. $(5c)^{\prime}$
    $K_{f}\equiv 6\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection of $f$. With the same arguments, we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+\frac{1}{2}+\frac{1}{2}+1\right)\mu_{1}>12\mu_{1}.$
  8. $(5d)^{\prime}$
    Let $K_{f}\equiv 6\Gamma+2\Gamma^{\prime}+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection. With the same arguments, we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+\frac{2}{3}+1\right)\mu_{1}>12\mu_{1}.$
  9. $(5e)^{\prime}$
    $K_{f}\equiv 6\Gamma+\Sigma+\Sigma^{\prime}+\mu_{1}F$ where $\Sigma,\Sigma^{\prime}$ are distinct bisections. With the same arguments, we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+1+1\right)\mu_{1}>12\mu_{1}.$
  10. $(5f)^{\prime}$
    $K_{f}\equiv 6\Gamma+3\Gamma+\Gamma^{\prime}+\mu_{1}F$ would imply
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{6}{7}+\frac{3}{4}+\frac{1}{2}\right)\mu_{1}=\left(10+\frac{28}{14}\right)\mu_{1}=12\mu_{1},$
    In this case we obtain a contradiction using point (4) of Proposition 6.4.
  11. $(5h)^{\prime}$
    $K_{f}\equiv 6\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a $4$-section. In this case we see in the very same way as above that $2K_{f}\cdot\Sigma\geq 4\mu_{1}$, so we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+2\right)\mu_{1}=12\mu_{1}.$
  12. $(6a)^{\prime}$
    Suppose that $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime}+\Gamma^{\prime\prime\prime\prime\prime}+\mu_{1}F$. By the same computations we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}\right)\mu_{1}>12\mu_{1}.$
  13. $(6b)^{\prime}$
    Suppose $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+2\Gamma^{\prime\prime\prime\prime}+\mu_{1}F$, We get
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{2}{3}\right)\mu_{1}>12\mu_{1}.$
  14. $(6c)^{\prime}$
    or $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Gamma^{\prime\prime}+\Gamma^{\prime\prime\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a bisection; then we get
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+1\right)\mu_{1}>12\mu_{1}.$
  15. $(6d)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+2\Gamma^{\prime\prime\prime}+\mu_{1}F$, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{2}{3}+\frac{2}{3}\right)\mu_{1}>12\mu_{1}.$
  16. $(6e)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+2\Gamma^{\prime\prime}+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{2}{3}+1\right)\mu_{1}>12\mu_{1}.$
  17. $(6g)^{\prime}$
    $K_{f}\equiv 5\Gamma+2\Gamma^{\prime}+3\Gamma^{\prime\prime}+\mu_{1}F$, we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{3}{4}\right)\mu_{1}>12\mu_{1}.$
  18. $(6f)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Gamma+\Sigma+\Sigma^{\prime}+\mu_{1}F$ where $\Sigma,\Sigma^{\prime}$ are distinct bisections, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+1+1\right)\mu_{1}>12\mu_{1}.$
  19. $(6h)^{\prime}$
    $K_{f}\equiv 5\Gamma+2\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a trisection, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{2}{3}+\frac{3}{2}\right)\mu_{1}>12\mu_{1}.$
  20. $(6i)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Sigma+3\Gamma^{\prime}+\mu_{1}F$, where $\Sigma$ is a bisection, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+1+\frac{3}{2}\right)\mu_{1}>12\mu_{1}.$
  21. $(6j)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+4\Gamma^{\prime\prime}+\mu_{1}F$, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+\frac{4}{2}\right)\mu_{1}>12\mu_{1}.$
  22. $(6k)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Gamma^{\prime}+\Sigma+\mu_{1}F$, where $\Sigma$ is a $4$-section, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{1}{2}+2\right)\mu_{1}>12\mu_{1}.$
  23. $(6m)^{\prime}$
    $K_{f}\equiv 5\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ is a $5$-section of $f$, then we obtain
    $K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq\left(10+\frac{5}{6}+\frac{5}{2}\right)\mu_{1}>12\mu_{1}.$
Remark 8.9.
As an example, let us consider one of the first cases that cannot be excluded by the above reasoning: case $(3c)$. Suppose indeed that $K_{f}\equiv 8\Gamma+\Sigma+\mu_{1}F$ where $\Sigma$ a bisection of $f$. As above, we can obtain
$9K_{f}\cdot\Gamma\geq\mu_{1}.$
and
$2K_{f}\cdot\Sigma\geq K_{f}\cdot\Sigma-\Sigma^{2}\geq 2\mu_{1}$
Now we compute $K_{f}^{2}$ as in the proof of the above theorem and obtain
$K_{f}^{2}=K_{f}\cdot Z_{1}+10\mu_{1}\geq 10\mu_{1}+8K_{f}\cdot\Gamma+K_{f}\cdot\Sigma\geq 10\mu_{1}+\frac{8}{9}\mu_{1}+\mu_{1},$
which does not give a contradiction with (2.11). Let us also consider the very last case: $K_{f}\equiv 5\Sigma+\mu_{1}F$ where $\Sigma$ is a bisection of $f$. Using the same computations, we get
$K_{f}^{2}\geq\left(10+\frac{10}{6}\right)\mu_{1},$
which does not give a contradiction.
So far we have excluded many numerical classes for $K_{f}$ just using Proposition 8.3 combined with the fact that the slope $K_{f}^{2}/\chi_{f}$ has to be smaller than $12$. Now we make the trigonal structure of $F$ come into play. We see that in the cases when $\alpha\geq 6$ we have a control on the possible Weierstrass sequence of the intersection of the section $\Gamma$ with the general fibre $F$.
Proposition 8.10.
In the situation of the theorem, if $\alpha>5$ then $p:=\Gamma\cdot F$ is a Weierstrass point for the curve. We have two possibilities:
  • (1)
    $p$ is a total ramification point for the trigonal morphism of $F$, whose non-gap sequence is one of the following:
    1. (i)
      (1, 2, 4, 5, 8, 11)1245811(1,\,2,\,4,\,5,\,8,\,11)( 1 , 2 , 4 , 5 , 8 , 11 ); this can only happen for $\alpha=10$;
    2. (ii)
      (1, 2, 4, 5, 7, 10)1245710(1,\,2,\,4,\,5,\,7,\,10)( 1 , 2 , 4 , 5 , 7 , 10 ); this can only happen for $\alpha=9$ or $\alpha=6$.
  • (2)
    $p$ is not a ramification point of the trigonal morphism of $F$, in this case the non-gap sequence of $p$ is as follows
    1. $(i)^{\prime}$
      (1, 2, 3, 4, 5, 11)1234511(1,\,2,\,3,\,4,\,5,\,11)( 1 , 2 , 3 , 4 , 5 , 11 ); this can only happen for $\alpha=10$;
    2. $(ii)^{\prime}$
      (1, 2, 3, 4, 5, 10)1234510(1,\,2,\,3,\,4,\,5,\,10)( 1 , 2 , 3 , 4 , 5 , 10 ); this can only happen for $\alpha=9$ or $\alpha=6$.
In particular the case (2) and (3) of Corollary 8.8 never can happen.
Proof.
First we prove that $p$ is a Weierstrass point of $F$ if $\alpha\geq 6$: indeed in this case we see that $h^{0}(6p)=h^{0}(K_{f}-R)$, where $R$ is an effective divisor on $F$ of degree $4$, so it holds $h^{0}(6p)\geq h^{0}(K_{F})-4=2$.
We want to determine non gap sequence of the point $p$.
Recall that we know that the fibre $F$ is trigonal with special Maroni invariant $n=2$ ([13]). Thanks to the work of Coppens [20, 21, 22] and Brundu-Sacchiero [14] we have a complete description of the possible non-gap sequences of a Weierstrass point on a trigonal curve. We list here the possibilities
  1. 1.
    [20, 22] if $p$ is a total ramification point of the trigonal morphism (i.e. $|3p|=g^{1}_{3}$), then the possible non-gap sequence are:
    (1, 2, 4, 5, 7, 10)1245710(1,\,2,\,4,\,5,\,7,\,10)( 1 , 2 , 4 , 5 , 7 , 10 ) (8.11)
    (1, 2, 4, 5, 8, 11)1245811(1,\,2,\,4,\,5,\,8,\,11)( 1 , 2 , 4 , 5 , 8 , 11 ) (8.12)
  2. 2.
    [14, Theorem 3.1] [22] If $p$ is a simple ramification point of of the trigonal morphism (i.e. $|2p+r|=g^{1}_{3}$, with $r\not=p$), then the possible non-gap sequence in $p$ are:
    (1, 2, 3, 4, 5, 7)123457(1,\,2,\,3,\,4,\,5,\,7)( 1 , 2 , 3 , 4 , 5 , 7 ) (8.13)
    (1, 2, 3, 4, 6, 8)123468(1,\,2,\,3,\,4,\,6,\,8)( 1 , 2 , 3 , 4 , 6 , 8 ) (8.14)
  3. 3.
    [14, Theorem 3.5] If $p$ is not a ramification point of of the trigonal morphism then there exist integers $r,\gamma$, with $4\leq r\leq 5$, $51\leq\gamma\leq 2r-$ 5 such that the possible non-gap sequence in $p$ are:
[…]
Let us start with case $\alpha=10$. We have that $10p\sim K_{F}$, so $h^{0}(11p)=h^{0}(K_{F}(p))=h^{0}(K_{F})=h^{0}(10p)=6$, hence the first non gap is $11$. So the only possibilities for the sequences are (8.12) or (8.15) with $r=\gamma=5$. The last sequence is
(1, 2, 3, 4, 5, 11).1234511(1,\,2,\,3,\,4,\,5,\,11).( 1 , 2 , 3 , 4 , 5 , 11 ) .
Let us now consider case $6\leq\alpha\leq 9$. So $K_{F}\sim\alpha p+R$, where $R$ is an effective divisor of degree $10-\alpha$ whose support does not contain $p$. Observe that $h^{0}(11p)=6$ by Riemann-Roch and $h^{0}(10p)=h^{0}(9p)=5$, because $10p\not\sim K_{F}$, and $9p\sim K_{F}-q$. So, the last non-gap number is $10$. By the results recalled above, the only possibilities are the sequences (8.11), or (8.15) with $r=5,\,\gamma=4$. In this last case the sequence is
(1, 2, 3, 4, 9, 10).1234910(1,\,2,\,3,\,4,\,9,\,10).( 1 , 2 , 3 , 4 , 9 , 10 ) .
Let us now see that this sequence and (8.11) can only happen for $\alpha=6$ or $9$. With both the sequences one sees immediately that $h^{0}(9p)=5$, so $h^{1}(9p)=1$, and this implies that $K_{F}\sim 9p+q$, where $q\not=p$.
Suppose that $K_{F}\sim 7p+R$ where $R$ is an effective divisor of degree $3$ not containing $p$. So, we have that $7p+R\sim 9p+q$, and so $R\sim 2p+q$. This would imply that $p$ is a simple ramification point for the trigonal map, which is impossible because we already ruled out this possibility.
Suppose now that $K_{F}\sim 8p+R$ where $R$ is an effective divisor of degree $2$ not containing $p$. With the same argument as above, we have $R\sim p+q$, so $F$ would be hyperelliptic, a contradiction. ∎
Remark 8.16.
With similar arguments one can prove that in case (6) of Corollary 8.8 if we consider the intersection $p=\Gamma\cap F$ or $p+q=\Sigma\cap F$, then $p$ need to be either a total ramification for the trigonal map or it a non ramified point.
We end this section with a similar result on genus five maximally irregular fibrations.
Proposition 8.17.
Let $f\colon S\to B$ a relatively minimal non locally trivial fibred surface of genus $g=5$ such that $u_{f}=4$. Then $f$ is a trigonal fibration, $S$ has positive index $\tau(S)>0$, and $Z_{1}$ has at least one multiple horizontal component, i.e. $\alpha\geq 2$.
Proof.
Let us assume that $Z_{1}$ does not have multiple vertical components. If $Z_{1}$ doesn’t have horizontal components, we know from Theorem 4.8 that we would have $K_{f}^{2}\geq 4\frac{g-1}{g-u_{f}}=16\chi_{f}$ which is impossible. If $Z_{1}$ is non-zero with reduced horizontal components we use (8.5) of Proposition 6.4 and obtain in any possible case a slope greater than $12$. Indeed, we have, give $Z_{1}=\sum_{j}G_{j}$ ($m_{j}=1$)
$K_{f}^{2}\geq\left(2(g-1)+\sum_{j=1}^{k}\frac{1}{2}b_{j}\right)\mu_{1}=\left(8+\frac{8}{2}\right)\mu_{1}=12\chi_{f},$
because $\sum_{j}b_{j}=2g-2=8$. So, we obtain a contradiction with point (3) of Theorem 5.2. ∎